### v1

# Community detection in heterogeneous signed networks

Abstract
Network data has attracted growing interest across scientific domains, prompting the development of various network models. Existing network analysis methods mainly focus on unsigned networks, whereas signed networks, consisting of both positive and negative edges, have been frequently encountered in practice but much less investigated. In this paper, we formally define strong and weak balance in signed networks, and propose a signed block $\beta$-model, which is capable of modeling strong- and weak-balanced signed networks simultaneously. We establish the identifiability of the proposed model by leveraging properties of bipartite graphs, and develop an efficient alternating updating algorithm to optimize the resulting log-likelihood function. More importantly, we establish the asymptotic consistencies of the proposed model in terms of both probability estimation and community detection. Its advantages are also demonstrated through extensive numerical experiments and the application to a real-world international relationship network.
Keywords: Balance theory, $\beta$-model, node heterogeneity, signed networks, stochastic block model
1 Introduction
Network data has attracted increasing attention in recent years, with applications in diverse scientific domains, ranging from social science [hunter2008goodness], computer science [heard2014filtering] to biomedicine [shojaie2009analysis]. Analysis of network data can reveal latent structures underlying the interactions among entities of interest, such as community detection [girvan2002community], link prediction [martinez2016survey], homophily [hoff2007modeling] and so on.
Most existing methods for network analysis focus on unsigned networks, whereas signed networks are also prevalent in real-world applications. Specifically, signed networks have been extensively employed to represent relationships among objects [harary1953notion], which associate polarity information with each edge and thus distinguish friends from enemies, or allies from rivals. These signed edges give rise to unique structural characteristics grounded in social psychology theories, particularly the balance theory [heider1946attitudes; cartwright1956structural], stating that certain triadic configurations involving signed edges lead to stable or unstable social structures. Figure 1 illustrates four possible triads, among which triads A and C represent strong-balanced structures. These triads align with intuitive social principles, such as “a friend of my friend is likely my friend” and “an enemy of my friend is likely my enemy”. Weak balance [davis1967clustering] extends this concept by additionally permitting triad D, which accommodates the occurrence of three mutually hostile nodes.
Figure 1: Configurations of four possible triads.
Community detection has been widely studied for unsigned networks [holland1983stochastic; girvan2002community; newman2006modularity; lei2015consistency; zhang2022directed], but much less so for signed networks. Most existing community detection methods for signed networks [doreian1996partitioning; kunegis2010spectral; jiang2015stochastic; cucuringu2021regularized] are algorithm based and lack formal theoretical analysis, not to mention that the prevalent balanced structures in signed networks are often ignored by the aforementioned community detection methods. Note that the balance theory actually posits that a strong-balanced (weak-balanced) network can be partitioned into two (multiple) sub-networks such that positive edges connect nodes within the same sub-network, whereas nodes from different sub-networks are connected by negative edges [easley2010networks]. More recently, tang2024population introduces a formal definition of strong-balanced signed networks in a probabilistic sense and proposes an efficient method to detect the two communities, yet the method is not readily applicable to the weak-balanced signed networks with multiple communities.
In this paper, we first extend the definition of strong-balanced signed networks in tang2024population to accommodate weak-balanced signed networks. We also provide a set of equivalence conditions for both strong- and weak-balanced signed networks, from both local and global perspectives. Built upon the global perspective, we then propose a signed block $\beta$-model to explicitly incorporate the strong or weak balance structures in heterogeneous signed networks. Particularly, it introduces two sets of parameters to capture each node’s distinct connection tendencies within its own community and across different communities. The identifiability of the signed block $\beta$-model is established by exploiting properties of the bipartite graph, and an efficient alternating updating algorithm is developed to optimize the resulting negative log-likelihood function for parameter estimation. More importantly, the asymptotic consistencies of the block $\beta$-model are established in terms of both probability estimation and community detection. Its advantages are also supported in numerical experiments on various synthetic signed networks and a real-world international relationship network.
The main contribution of the proposed signed block $\beta$-model is that it provides a novel framework to statistically model strong and weak balance structures in signed networks. It also generalizes the traditional $\beta$-model [chatterjee2011random; 10.1214/12-AOS1078; chen2021analysis] by incorporating community structures while preserving its capacity to capture node heterogeneity. For each node, two sets of parameters are introduced to model its tendency to positively or negatively connect with other nodes within the same community or across different communities. Such a treatment not only incorporates community structure in each node’s connection tendency, but also admits more substantial node heterogeneity than existing models [chen2021analysis; shao20212; cucuringu2021regularized]. Moreover, a novel set of identifiability conditions for the proposed signed block $\beta$-model is established by leveraging techniques from the graph theory. The theoretical techniques developed here can be applicable to identifiability analyses of other network models involving community structures, and thus are of independent interest.
The remainder of this paper is organized as follows. Section 2 formally gives the definition of strong and weak balance in signed networks. Section 3 presents the signed block $\beta$-model, establishes its identifiability conditions, and develops an efficient alternative updating scheme to tackle the resulting optimization task. Section 4 provides a rigorous theoretical analysis of the proposed signed block $\beta$-model and quantifies its asymptotic behaviors in both probability estimation and community detection. Section 5 examines the numeric performance of the signed block $\beta$-model via extensive experiments on both synthetic and real-world signed networks. Finally, Section 6 concludes the paper with a brief summary. Technical proofs of the main theorems are provided in the Appendix, whereas auxiliary lemmas are contained in the Supplementary Material.
Notations.
We introduce some notations to be used throughout the paper. For any integer $n$, let $n]=\{1,\dots,n\}$. For any vector $\bm{v}=(v_{1},\dots,v_{n})^{\top}\in\mathcal{R}^{n}$, $\|\bm{v}\|_{2}^{2}=\sum_{i\in[n]}v_{i}^{2}$, and $\|\bm{v}\|_{\infty}=\max_{i\in[n]}|v_{i}|$. We use $\bm{1}_{n}$ to denote the $n$-dimensional column vector with all elements being 1 and $\mathbb{I}(\cdot)$ to denote the indicator function. For any matrix $\bm{A}=(a_{ij})\in\mathcal{R}^{n\times m}$, let $\bm{A}_{i}$ be its $i$-th row, and $\|\bm{A}\|_{F}^{2}=\sum_{i\in[n],j\in[m]}a_{ij}^{2}$ be its Frobenius norm. For any two matrices $\bm{A}$ and $\bm{B}$, $\bm{A}\odot\bm{B}$ is the Hadamard product with $(\bm{A}\odot\bm{B})_{ij}=a_{ij}b_{ij}$. For any two sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we denote $a_{n}\prec b_{n}$ if $a_{n}=o(b_{n})$, $a_{n}\preceq b_{n}$ if $a_{n}=O(b_{n})$ and $a_{n}\simeq b_{n}$ if $a_{n}\preceq b_{n}$ and $b_{n}\preceq a_{n}$. For any set $\mathcal{X}$, $|\mathcal{X}|$ denotes its cardinality and $\mathcal{X}^{c}$ denotes its complement. For any two sets $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{X}\setminus\mathcal{Y}=\{x\mid x\in\mathcal{X}\text{ and }x\notin\mathcal{Y}\}$.
2 Signed network and balance theory
A signed network $\mathcal{G}$ can be represented via its symmetric adjacency matrix $\bm{A}=(a_{ij})_{i,j\in[n]}$ with $a_{ij}\in\{1,-1,0\}$, where $a_{ij}=1$ indicates a positive edge between nodes $i$ and $j$, $a_{ij}=-1$ indicates a negative edge and $a_{ij}=0$ indicates no edge at all. For each $i,j\in[n]$, let $a_{ij}$ be sampled from a multinomial distribution with parameters $(p_{ij}^{+},p_{ij}^{-},p_{ij}^{0})$. One of the most salient features of signed network is the widely observed balance structure [harary1953notion; davis1967clustering], positing that relationships among entities tend toward balanced sign configurations.
Definition 1 (Strong balance)
A signed network is strong-balanced if for any triple $(i,j,k)$,
$\underset{(a,b,c)\in\{-1,1\}^{3}}{\arg\max}\mathbb{P}\big(a_{ij}=a,a_{jk}=b,a_{ki}=c\big)$
does not have 0 or 2 positive edges.
Definition 1 assures that a strong-balanced signed network does not contain a triad achieving the highest probability at triad B or D as shown in Figure 1. An equivalent definition has also been introduced in tang2024population, which quantifies the probabilistic strong balance for signed networks. Moreover, Definition 1 can be extended to define weak-balanced signed networks.
Definition 2 (Weak balance)
A signed network is weak-balanced if for any triple $(i,j,k)$,
$\underset{(a,b,c)\in\{-1,1\}^{3}}{\arg\max}\mathbb{P}\big(a_{ij}=a,a_{jk}=b,a_{ki}=c\big)$
does not have exactly 2 positive edges.
Definition 2 gives a probabilistic definition of weak balance for signed networks. It only restricts the occurrence of triad B, and thus aligns with the non-probabilistic definitions of weak balance in the existing literature [easley2010networks].
Proposition 1
A signed network $\mathcal{G}$ is strong-balanced if and only if
(i) For any $(i,j,k)$, $(p_{ij}^{+}-p_{ij}^{-})(p_{jk}^{+}-p_{jk}^{-})(p_{ki}^{+}-p_{ki}^{-})>0$; or
(ii) There exists a partition $n]=C_{1}\cup C_{2}$ such that for each $i,j\in[n]$, $p_{ij}^{+}>p_{ij}^{-}$ if $i,j$ belong to the same subset, and $p_{ij}^{+}<p_{ij}^{-}$, if $i,j$ belong to different subsets.
Proposition 2
A signed network $\mathcal{G}$ is weak-balanced if and only if
(i) For any $(i,j,k)$, $(p_{ij}^{+}-p_{ij}^{-})(p_{jk}^{+}-p_{jk}^{-})(p_{ki}^{+}-p_{ki}^{-})>0$ or $p_{ij}^{+}-p_{ij}^{-}<0,p_{jk}^{+}-p_{jk}^{-}<0,p_{ki}^{+}-p_{ki}^{-}<0$; or
(ii) There exists a partition of $n]=C_{1}\cup C_{2}\cup...\cup C_{K}$ with $K\geq 3$ such that for each $i,j\in[n]$, $p_{ij}^{+}>p_{ij}^{-}$ if $i,j$ belong to the same subset, and $p_{ij}^{+}<p_{ij}^{-}$, if $i,j$ belong to different subsets.
Propositions 1 and 2 characterize the strong and weak balance structures in signed networks from both local and global perspectives. In fact, the local perspective (i) in Proposition 1 is essentially equivalent to $\mathbb{E}(a_{ij}a_{jk}a_{ki}\big||a_{ij}a_{jk}a_{ki}|=1)>0$ for any $(i,j,k)$, which is exactly the definition of the strong-balanced signed network in tang2024population, whereas (i) in Proposition 2 can be regarded as its analogy for the weak-balanced signed network. On the other hand, (ii) in Proposition 1 and (ii) in Proposition 2 give the global perspectives on strong and weak balance structures, which greatly facilitate the modeling of community structures in signed networks.
3 Signed block $\beta$-model
We propose a signed block $\beta$-model, denoted as SBBM, to leverage the global perspectives of the balance structures in signed networks. Particularly, let $\psi_{i}\in[K]$ denote the community membership of node $i$, then
[…]
for each $1\leq i\leq j\leq n$, and $\mathbb{P}(a_{ij}=0)=1-\mathbb{P}(a_{ij}=1)-\mathbb{P}(a_{ij}=-1)$. It is clear that SBBM associates each node $i$ with two sets of parameters $(\beta_{i}^{+},\beta_{i}^{-})$ and $(\gamma_{i}^{+},\gamma_{i}^{-})$, where $(\beta_{i}^{+},\beta_{i}^{-})$ quantifies the node’s tendency to form positive and negative edges within the same community, and $(\gamma_{i}^{+},\gamma_{i}^{-})$ characterizes the corresponding tendencies across different communities.
The proposed SBBM model has a number of compelling features. First, it integrates the capability of community detection while simultaneously capturing node-specific heterogeneity in signed networks. When $K=1$ and $\beta_{i}^{-}=$-$\infty$, SBBM reduces to the classical $\beta$-model, which is commonly used to capture node-specific heterogeneity. In the absence of negative edges, SBBM simplifies to the stochastic block model (SBM) with heterogeneity for community detection. Therefore, SBBM provides a general framework for both unsigned and signed networks, simultaneously addressing community detection, node-specific heterogeneity and signed relationships in network data. Second, both strong and weak balance can be incorporated in SBBM by simply requiring
$\beta_{i}^{+}>\beta_{i}^{-}\mbox{ and }\gamma_{i}^{+}<\gamma_{i}^{-}\ \text{ for any }i\in[n].$ (2)
It can be readily verified that nodes $i$ and $j$ belong to the same community if and only if $\mathbb{P}(a_{ij}=1)>\mathbb{P}(a_{ij}=-1)$, and different communities if and only if $\mathbb{P}(a_{ij}=1)<\mathbb{P}(a_{ij}=-1)$. It follows from Propositions 1 and 2 that SBBM with constraints (2) satisfies strong balance when $K=2$ and weak balance when $K\geq 3$.
3.1 Model identifiability
Let $\bm{Z}=(\bm{z}_{1},\dots,\bm{z}_{n})^{\top}$ be the community membership matrix, where $\bm{z}_{i}\in\{0,1\}^{K}$ with $z_{ik}=1$ if $\psi_{i}=k$. Define $\bm{\beta}^{+}=(\beta^{+}_{i})_{i=1}^{n}$, $\bm{\gamma}^{+}=(\gamma_{i}^{+})_{i=1}^{n}$, $\bm{\beta}^{-}=(\beta^{-}_{i})_{i=1}^{n}$, $\bm{\gamma^{-}}=(\gamma_{i}^{-})_{i=1}^{n}$, $\bm{\eta}^{+}=\bm{\beta}^{+}-\bm{\gamma}^{+}$ and $\bm{\eta}^{-}=\bm{\beta}^{-}-\bm{\gamma}^{-}$. Then, the parameter space is
$\mathcal{P}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ $\times\mathbb{R}^{n}\times\{0,1\}^{n\times n}:$
[…]
where $\bm{Z}\bm{1}_{K}=\bm{1}_{n}$ assures that each node belongs to exactly one community. Moreover, let $\bm{\Theta}^{+}=(\theta_{ij}^{+})_{i,j\in[n]}$ and $\bm{\Theta}^{-}=(\theta_{ij}^{-})_{i,j\in[n]}$, where $\theta_{ij}^{+}=\log\frac{\mathbb{P}(a_{ij}=1)}{1-{\mathbb{P}(a_{ij}=1)-\mathbb{P}(a_{ij}=-1)}}$ and $\theta_{ij}^{-}=\log\frac{\mathbb{P}(a_{ij}=-1)}{1-\mathbb{P}(a_{ij}=1)-{\mathbb{P}(a_{ij}=-1)}}$. Then, SBBM in (1) can be rewritten as
$\begin{cases}\bm{\Theta}^{+}=\bm{\gamma}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\bm{\gamma}^{+\top}+(\bm{\eta}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\bm{\eta}^{+\top})\odot\bm{Z}\bm{Z}^{\top},\\ \bm{\Theta}^{-}=\bm{\gamma}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\bm{\gamma}^{-\top}+(\bm{\eta}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\bm{\eta}^{-\top})\odot\bm{Z}\bm{Z}^{\top}.\end{cases}$
We say SBBM is identifiable if parameters $(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}\bm{Z}^{\top})$ can be uniquely determined from $\bm{\Theta}^{+}$ and $\bm{\Theta}^{-}$, where $\bm{Z}$ can be uniquely determined up to column permutation.
Theorem 1
SBBM is identifiable when $K=2$, if $\eta_{1}^{+}=0$, $\eta_{1}^{-}=0$ and there does not exist $S_{1}\subseteq C_{1}$ and $S_{2}\subseteq C_{2}$ such that $\eta_{i}^{+}+\eta_{j}^{+}=0$ or $\eta_{i}^{-}+\eta_{j}^{-}=0$ for any $i\in S_{1},j\in S_{2}$ and $i\in C_{1}\setminus S_{1},j\in C_{2}\setminus S_{2}$.
SBBM is also identifiable when $K\geq 3$, if
(i) there does not exist a community $C$ and a subset $S\subseteq C$ such that $\eta_{i}^{+}+\eta_{j}^{+}=0$ or $\eta_{i}^{-}+\eta_{j}^{-}=0$ for any $i\in S$ and $j\in C\setminus S$; and
(ii) there does not exist two communities $C$ and $C^{\prime}$ such that $\eta_{i}^{+}+\eta_{j}^{+}=0$ or $\eta_{i}^{-}+\eta_{j}^{-}=0$ for any $i\in C$ and $j\in C^{\prime}$.
Intuitively, SBBM is unidentifiable if the difference between $\bm{\beta}^{+}$ and $\bm{\gamma}^{+}$, or between $\bm{\beta}^{-}$ and $\bm{\gamma}^{-}$ is not substantial, especially when their effects intertwine with the community membership. Theorem 1 establishes the identifiability of SBBM by assuring that there are not too many node pairs with $\eta_{i}^{+}+\eta_{j}^{+}=0$ or $\eta_{i}^{-}+\eta_{j}^{-}=0$ so that the connection probability within the same community and across different communities can be well discriminated [latouche2011overlapping; zhang2022identifiability]. The case with $K=2$ requires separate treatment, as the underlying network distribution remains invariant when all $\eta^{+}$’s or $\eta^{-}$’s in one community increase but decrease in the other. Proof of Theorem 1 involves a detailed graph-theoretic analysis of the intertwined relationship between the parameters and community membership, which may be of independent interest.
3.2 Computing algorithm
The negative log-likelihood function of SBBM is
$L(\bm{\Theta};{\bm{A}})=-\frac{2}{n(n+1)}\sum_{1\leq i\leq j\leq n}\Big(a^{+}_{ij}\theta^{+}_{ij}+a^{-}_{ij}\theta^{-}_{ij}-\log\big(1+e^{\theta^{+}_{ij}}+e^{\theta^{-}_{ij}}\big)\Big),$ (3)
where $\bm{\Theta}=(\bm{\Theta}^{+},\bm{\Theta}^{-})$, $a_{ij}^{+}=\mathbb{I}(a_{ij}=1)$, $a_{ij}^{-}=\mathbb{I}(a_{ij}=-1)$ for each $1\leq i\leq j\leq n$. Define
$\mathcal{P}_{0}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathcal{P}:\$$\eta_{i}^{+}+\eta_{j}^{+}=0\ \text{or}\ \eta_{i}^{-}+\eta_{j}^{-}=0$
$\text{for any}\ i\in S_{1},j\in S_{2}\ \text{and}\ i\in S_{1}^{c},j\in S_{2}^{c}\big\},$
where $S_{1}\subseteq C_{1}$ and $S_{2}\subseteq C_{2}$, and
[…]
Then, $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ can be obtained by minimizing $L(\bm{\Theta};{\bm{A}})$ over $(\mathcal{P}\setminus\mathcal{P}_{0})\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$ and over $\mathcal{P}\setminus\mathcal{P}_{1}$ when $K\geq 3$. It is clear that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ are zero sets under the Lebesgue measure. Therefore, we get $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ by minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$, and minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}$ directly when $K\geq 3$ [latouche2011overlapping; zhang2022identifiability].
We develop an alternative updating scheme to optimize $L(\bm{\Theta};{\bm{A}})$. We first rewrite $L(\bm{\Theta};\bm{A})$ as $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})$ and initialize $\bm{\gamma}^{+(0)}=\bm{\eta}^{+(0)}=\bm{\gamma}^{-(0)}=\bm{\eta}^{-(0)}=\bm{0}_{n}$ and $\bm{Z}^{(0)}$ by applying spectral clustering algorithm to $\bm{A}$. Given $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})$ at the tt‑th step for $t\geq 1$, note that, with $\bm{Z}^{(t-1)}$ fixed, minimizing $L(\bm{\Theta};\bm{A})$ subject to (2) with respect to $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)})$ is a convex optimization problem. We therefore adopt the Spectral Projected Gradient method [birgin2000nonmonotone], which combines a Barzilai–Borwein spectral step, Euclidean projection onto the feasible set, and an Armijo back-tracking line search, thereby handling curvature adaptation and step-size control automatically. Upon convergence, we obtain the updated parameters $(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)})$. Note that we project $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ such that $\eta_{1}^{+(t)}=\eta_{1}^{-(t)}=0$ when $K=2$, which are still denoted as $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ with a slight abuse of notation.
Next, we adopt the similar method as in yuan2021community and zhang2024consistent to update $\bm{Z}^{(t)}$ due to its discrete nature. Let $\bm{Z}_{-i}$ be a matrix by deleting of the ii‑th row of $\bm{Z}$. Then, node $i$ will be more likely to belong to the kk‑th community if $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{ik}=1)\leq L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{il}=1)$ for each $l\neq k$. Thus, we have
$z_{ik}^{(t)}=\begin{cases}1,&k=\displaystyle\arg\min_{l\in[K]}L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}_{-i}^{(t-1)},z_{il}=1\right);\\ 0,&\text{otherwise},\end{cases}$
for each $i\in[n]$ and $k\in[K]$.
After updating $\bm{Z}^{(t)}$, we check whether
$L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)}\right)<L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t-1)}\right).$
Otherwise, we update each $\bm{z_{i}}$ sequentially, thereby ensuring that the negative log‑likelihood always decreases. Let $\alpha$ and $T_{\max}$ be threshold values for the relative error and updating steps. The whole algorithm will stop when $t>T_{\max}$ or the relative error
$\frac{\bigl|L(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)})-L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}{\bigl|L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}<\alpha.$
Let $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ be the output of the alternative updating scheme, then $\widehat{\bm{\Theta}}^{+}=\widehat{\bm{\gamma}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{+\top}+(\widehat{\bm{\eta}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{+\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$ and $\widehat{\bm{\Theta}}^{-}=\widehat{\bm{\gamma}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{-\top}+(\widehat{\bm{\eta}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{-\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$.
4 Theory
In this section, we establish the asymptotic properties of the proposed SBBM in terms of both probability estimation and community detection. Denote the true parameters as $(\bm{\beta}^{*+},\bm{\beta}^{*-},\bm{\gamma}^{*+},\bm{\gamma}^{*-},\bm{Z}^{*})\in\mathcal{P}$ with $\bm{\beta}^{*+}>\bm{\beta}^{*-}$ and $\bm{\gamma}^{*+}<\bm{\gamma}^{*-}$, and $\bm{\Theta}^{*}=(\bm{\Theta}^{*+},\bm{\Theta}^{*-})$ correspondingly.
4.1 Consistency in probability estimation
For each $i,j\in[n]$, denote the negative log-likelihood of $a_{ij}$ as
$L(\theta_{ij};a_{ij})=-(a_{ij}^{+}\theta_{ij}^{+}+a_{ij}^{-}\theta_{ij}^{-})+\log\big(1+e^{\theta_{ij}^{+}}+e^{\theta_{ij}^{-}}\big),$
where $\theta_{ij}=(\theta_{ij}^{+},\theta_{ij}^{-})$, and $L(\bm{\Theta};\bm{A})=\frac{2}{n(n+1)}\sum_{i\leq j}L(\theta_{ij};a_{ij})$. Let $e_{L}(\bm{\Theta},\bm{\Theta}^{*})=\mathbb{E}\big(L(\bm{\Theta};\bm{A})-L(\bm{\Theta}^{*};\bm{A})\big)$, which serves as an intermediate error metric measuring the accuracy of $\bm{\Theta}$ in estimating $\bm{\Theta}^{*}$.
Assumption 1
There exist constants $M_{1}<M_{2}$ such that both $\theta_{ij}^{+}$ and $\theta_{ij}^{-}$ are contained in $M_{1},M_{2}]$ for each $i,j\in[n]$.
Assumption 1 is common in the literature of network modeling. It allows different magnitudes of $M_{1}$ and $M_{2}$, leading to highly heterogeneous connecting probabilities. This is in sharp contrast to the homogeneous assumption on connecting probabilities in most existing literature on sparse networks, where $M_{1}$ and $M_{2}$ are essentially of the same magnitude [zhang2022directed; li2023statistical; lei2023bias; zhen2023community].
Proposition 3
For any $\bm{\Theta}$ that satisfies Assumption 1, it holds that
$\frac{2}{n(n+1)}\sum_{i\leq j}\mathbb{E}[L(\theta_{ij},a_{ij})-L(\theta^{*}_{ij},a_{ij})]^{2}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}e_{L}(\bm{\Theta},\bm{\Theta}^{*}),$
$\frac{1}{n^{2}}\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}e_{L}(\bm{\Theta},\bm{\Theta}^{*}).$
Proposition 3 paves a bridge between the first and second moments of $L(\bm{\Theta},\bm{A})-L(\bm{\Theta}^{*},\bm{A})$, which is crucial for establishing the weak consistency of $\widehat{\bm{\Theta}}$ in terms of $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$. Furthermore, it also relates $\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ to $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$, leading to the strong consistency of $\widehat{\bm{\Theta}}$ established in the following Theorem 2.
Theorem 2
[…]
$\ P\Big(\frac{1}{n^{2}}\big\|\widehat{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}\geq 48e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\Big)\leq 2\exp(-c_{1}Kn\log n),$
when $n$ is large enough, provided that $\epsilon_{n}^{2}\simeq\frac{Ke^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}\log n}{n}\prec 1$.
Theorem 2 guarantees that $\bm{\Theta}^{*}$ can be consistently estimated if $K\simeq 1$ and $\min\{2e^{M_{2}}+e^{2M_{2}},1\}e^{2\max\{|M_{1}|,|M_{2}|\}}n^{-1}\log n$ is vanishing. In particular, if $M_{2}-M_{1}\preceq 1$ and both of them diverge to -$\infty$, then it is required that $\min\{p_{ij}^{*+},p_{ij}^{*-}\}\succ\frac{\log n}{n}$ for each $i,j\in[n]$, which matches the best sparsity requirement in the literature [lei2015consistency; ma2020universal; zhang2022directed]. More importantly, when $M_{2}-M_{1}\succ 1$, $p_{ij}^{*+}$ and $p_{ij}^{*-}$ can behave differently to accommodate edge heterogeneity. Specifically, letting $M_{1}\to$-$\infty$ and $M_{2}\to\infty$, Theorem 2 still holds when $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$, resulting in a heterogeneous case with $p_{ij}^{*+},p_{ij}^{*-}\in(n^{-1/2}(\log n)^{1/2},1)$ for each $i,j\in[n]$. This is consistent with the heterogeneous result for the latent space model [ma2020universal; zhang2022joint], and better than that for the classical $\beta$-model [shao20212], where the connecting probabilities can only vary in $(n^{-1/4},1)$.
4.2 Consistency in community detection
We next turn to establish the asymptotic properties of $\widehat{\bm{Z}}$, whose community detection error is defined as
$d(\widehat{\bm{Z}},\bm{Z}^{*})=\min_{\bm{\Pi}}\frac{1}{n}\|\widehat{\bm{Z}}\bm{\Pi}-\bm{Z}^{*}\|_{F}^{2},$ (4)
where minimization is taken over all permutation matrices $\bm{\Pi}\in\{0,1\}^{K\times K}$. The quantity $d(\widehat{\bm{Z}},\bm{Z}^{*})$ measures the minimal fraction of misclassified nodes, up to permutation, and is commonly used in the community detection literature [lei2015consistency].
Assumption 2
There exists a constant $M_{\gamma}$ such that $\|\bm{\gamma}\|_{\infty}\leq M_{\gamma}$. Let $\mathcal{S}_{i}^{*+}=\big\{j\in[n]:\ |\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}\big\}$ and $\mathcal{S}_{i}^{*-}=\big\{j\in[n]:\ |\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}\big\}$ for some $c_{\eta}>0$, then $\max_{i\in[n]}|\mathcal{S}_{i}^{*+}|\prec\frac{n}{K}$ or $\max_{i\in[n]}|\mathcal{S}_{i}^{*-}|\prec\frac{n}{K}$.
Assumption 2 assures that the number of pairs of $(i,j)$ satisfying $|\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}$ or $|\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}$ is substantially less than $n/K$. Note that $|\beta_{i}^{*+}+\beta_{j}^{*+}|\geq 2M_{\gamma}+c_{\eta}$ and $|\beta_{i}^{*-}+\beta_{j}^{*-}|\geq 2M_{\gamma}+c_{\eta}$ imply that $\min\{|\eta_{i}^{*+}+\eta_{j}^{*+}|,|\eta_{i}^{*-}+\eta_{j}^{*-}|\}\geq c_{\eta}$, thus $c_{\eta}$ can be regarded as a measurement of the difference in the level of connection probability within the same community and those across different communities. For any $(\bm{\gamma}^{+},\bm{\gamma}^{-},\bm{\eta}^{+},,\bm{\eta}^{-},\bm{Z})\in\mathcal{P}$, let $\mathcal{N}_{k}=\{i\in[n]:z_{ik}=1\}$ and $n_{k}=|\mathcal{N}_{k}|$ for each $k\in[K]$.
Proposition 4
Suppose all assumptions in Theorem 2 and Assumption 2 hold, and that $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$. Then, with probability at least $1-2\exp(-c_{1}Kn\log n)$, there does not exist $k_{1}\neq k_{2}\in[K]$ and $k\in[K]$ such that $|\mathcal{N}_{k_{1}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq|\mathcal{N}_{k_{2}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq\frac{n}{K}$.
The requirement of $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$ implies that $c_{\eta}$ can not be too small, which can be interpreted as a minimal signal strength condition. Proposition 4 assures that any estimated community $\widehat{\mathcal{N}}_{k}$ does not substantially overlap with more than one true community with high probability, which further leads to the estimation consistency of $\widehat{\bm{Z}}$ as shown in Theorem 3.
Theorem 3
Suppose all assumptions in Proposition 4 hold, then there exists a positive constant $c_{2}$ such that when $n$ is sufficiently large,
$P\big(d(\widehat{\bm{Z}},\bm{Z}^{*})\geq c_{2}Kc_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\big)\leq 2\exp(-c_{1}Kn\log n).$
Theorem 3 establishes the convergence rate of $d(\widehat{\bm{Z}},\bm{Z}^{*})$ for community detection. Particularly, in the homogeneous case with $\max\{|M_{1}|,|M_{2}|,K,c_{\eta}\}\simeq 1$, the convergence rate becomes $n^{-1}\log n$, which matches with the best community detection result in literature [lei2015consistency; zhang2022directed]. In the heterogeneous case with $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$ where $p_{ij}^{*}$ is allowed to vary in $(n^{-1/2}(\log n)^{1/2},1)$ for each $i\leq j$, $c_{\eta}$ can diverge to infinity at a rate no faster than $\log n$ and Theorem 3 continues to hold with $d(\widehat{\bm{Z}},\bm{Z}^{*})=O_{p}\big((\log n)^{-2}\big)$.
5 Numerical experiments
In this section, we examine the numerical performance of the proposed SBBM model, and compare it against three popular competitors in the literature, including the signed Laplacian-based method [kunegis2010spectral], the SPONGE algorithm based on signed SBM [cucuringu2021regularized], and the joint estimation method based on the joint inner product model [tang2024population], denoted as SLP, SPONGE and JIM, respectively. All methods are evaluated in terms of community detection accuracy, while JIM and SBBM are further compared in terms of probability estimation.
Note that the community detection error in (4) requires enumeration of all possible permutations, which is computationally infeasible. We thus employ the clustering error in wang2010consistent to evaluate the community detection accuracy. Particularly, let $\bm{\psi}^{*}=(\psi_{1}^{*},\dots,\psi_{n}^{*})$ and $\widehat{\bm{\psi}}=(\widehat{\psi}_{1},\dots,\widehat{\psi}_{n})$ denote the true and estimated community assignments, then the community detection error of $\widehat{\bm{\psi}}$ can be evaluated as
$\text{Err}(\widehat{\bm{\psi}},\bm{\psi}^{*})=\frac{2}{n(n-1)}\sum_{i<j}\mathbb{I}\big(\mathbb{I}(\psi_{i}^{*}=\psi_{j}^{*})+\mathbb{I}(\widehat{\psi}_{i}=\widehat{\psi}_{j})=1\big).$
Further, we define the estimation error of the probability matrix $\bm{P}^{*+}=(p^{*+}_{ij})_{i,j=1}^{n}$ and $\bm{P}^{*-}=(p^{*-}_{ij})_{i,j=1}^{n}$ as
[…]
5.1 Simulated examples
Three simulated signed networks are considered.
Example 1. Following the data generating scheme in tang2024population, we first generate $\bm{X}\in\mathbb{R}^{n\times 2}$ with rows independently set to $(1,0)$ or $(0,1)$ with equal probability $0.5$, and then set $\bm{X}^{*}=(\bm{I}_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\top})\bm{X}$. For each node $i\in[n]$, we generate $\alpha_{i}^{*}\sim\mathrm{Unif}(1,3)$ and let $v_{i}^{*}=(1,-1)^{\top}\bm{x}_{i}^{*}/\sqrt{2}$, where $\bm{x}_{i}^{*}$ denotes the $i$-th row of $\bm{X}^{*}$. Edge probabilities are then specified by $\Pr(|A_{ij}|=1)=\sigma(\mu^{*}+\alpha_{i}^{*}+\alpha_{j}^{*}-2\bar{\alpha}^{*}+\bm{x}_{i}^{*\top}\bm{x}_{j}^{*})$ and $\Pr(A_{ij}=1\mid|A_{ij}|=1)=\sigma(v_{i}^{*}v_{j}^{*})$, where $\bar{\alpha}^{*}=n^{-1}\sum_{i=1}^{n}\alpha_{i}^{*}$ and $\sigma(x)=(1+e^{-x})^{-1}$. Varying $\mu^{*}\in\{-3.5,-3.0,-2.5\}$ produces strong-balanced signed networks with different sparsity.
Example 2. In this example, for each $i\in[n]$, we first generate $\beta_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, $\gamma_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$, where $\delta_{i}\sim U(0,2)$. Then we vary $K\in\{4,6,8\}$ and randomly assign each node to one of the $K$ communities uniformly. The matrices $\bm{\Theta}^{*}$ and $\bm{P}^{*}$ are constructed as in (1) and the signed networks are generated accordingly, resulting in weak-balanced signed networks with different numbers of communities.
Example 3. In this example, we set $K=4$. For each $i\in[n]$, we generate $\beta_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$, $\gamma_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$ with $\sigma^{2}\in\{1,2,3\}$. We then generate $\delta_{i}\sim U(0,2)$ and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$. The signed networks are subsequently generated as in Example 2, resulting in a weak-balanced network with various node heterogeneity.
We consider $n\in\{500,1000\}$ in all examples. The averaged probability estimation errors and community detection errors, along with their standard errors, over 50 independent replications are reported in Tables 1-6.
Table 1: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 1. The best performer in each scenario is boldfaced.
$n$ $\mu$ SBBM SLP SPONGE JIM
500 -3.5 0.0910(0.0122) 0.4172(0.0069) 0.3950(0.0106) 0.2073(0.0242)
-3.0 0.0271(0.0014) 0.2822(0.0059) 0.2164(0.0049) 0.1060(0.0189)
-2.5 0.0078(0.0007) 0.1715(0.0050) 0.1169(0.0034) 0.0361(0.0106)
1000 -3.5 0.0133(0.0007) 0.2016(0.0029) 0.1515(0.0029) 0.0723(0.0162)
-3.0 0.0025(0.0003) 0.1070(0.0023) 0.0711(0.0018) 0.0230(0.0058)
-2.5 0.0004(0.0001) 0.0440(0.0012) 0.0257(0.0009) 0.0161(0.0100)
Table 2: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 1. The better performer in each scenario is boldfaced.
$n$ $\mu$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 -3.5 0.6085(0.1521) 0.8727(0.2557) 0.5546(0.0126) 0.4829(0.0106)
-3.0 0.3136(0.0014) 0.3964(0.0017) 0.4163(0.0091) 0.3848(0.0094)
-2.5 0.2582(0.0012) 0.3199(0.0011) 0.3187(0.0067) 0.3033(0.0079)
1000 -3.5 0.2789(0.0008) 0.3527(0.0010) 0.3726(0.0063) 0.3528(0.0069)
-3.0 0.2277(0.0007) 0.2850(0.0007) 0.2819(0.0032) 0.2805(0.0067)
-2.5 0.1888(0.0005) 0.2350(0.0005) 0.2170(0.0048) 0.2140(0.0065)
Table 3: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 2. The best performer in each scenario is boldfaced.
$n$ $K$ SBBM SLP SPONGE JIM
500 4 0.0310(0.0041) 0.3201(0.0063) 0.2276(0.0073) -
6 0.0425(0.0036) 0.3621(0.0049) 0.2692(0.0050) -
8 0.0508(0.0039) 0.3800(0.0056) 0.2886(0.0051) -
1000 4 0.0091(0.0006) 0.3128(0.0050) 0.2141(0.0060) -
6 0.0123(0.0011) 0.3497(0.0037) 0.2600(0.0038) -
8 0.0152(0.0012) 0.3632(0.0031) 0.2736(0.0032) -
Table 4: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 2. The better performer in each scenario is boldfaced.
$n$ $K$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 4 0.2265(0.0083) 0.1945(0.0049) 1.0216(0.0093) 0.8595(0.0035)
6 0.2841(0.0119) 0.1826(0.0038) 1.1016(0.0123) 0.8305(0.0033)
8 0.3409(0.0128) 0.1819(0.0044) 1.1649(0.0149) 0.8164(0.0033)
1000 4 0.1546(0.0035) 0.1344(0.0024) 1.0343(0.0096) 0.8587(0.0026)
6 0.1788(0.0049) 0.1225(0.0020) 1.1188(0.0130) 0.8290(0.0025)
8 0.1993(0.0053) 0.1168(0.0016) 1.1869(0.0159) 0.8151(0.0024)
Table 5: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 3. The best performer in each scenario is boldfaced.
$n$ $\sigma^{2}$ SBBM SLP SPONGE JIM
500 1 0.1858(0.0063) 0.2355(0.0071) 0.2603(0.0229) -
2 0.1418(0.0082) 0.2802(0.0073) 0.1675(0.0089) -
3 0.0784(0.0073) 0.2951(0.0051) 0.1662(0.0080) -
1000 1 0.0845(0.0032) 0.1855(0.0074) 0.0970(0.0077) -
2 0.0465(0.0015) 0.2640(0.0044) 0.0882(0.0049) -
3 0.0226(0.0010) 0.2749(0.0044) 0.1208(0.0080) -
Table 6: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 3. The better performer in each scenario is boldfaced.
$n$ $\sigma^{2}$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 1 0.4052(0.0063) 0.8484(0.0184) 1.6358(0.0486) 3.8576(0.0938)
2 0.3283(0.0089) 0.4778(0.0203) 1.2363(0.0253) 1.5960(0.0327)
3 0.2717(0.0092) 0.3165(0.0103) 1.1129(0.0168) 1.1200(0.0138)
1000 1 0.2725(0.0026) 0.5233(0.0055) 1.6849(0.0516) 3.7927(0.0585)
2 0.2084(0.0021) 0.2893(0.0035) 1.2639(0.0263) 1.5614(0.0211)
3 0.1751(0.0021) 0.2085(0.0024) 1.1328(0.0175) 1.1076(0.0092)
It is evident that SBBM delivers the best performance in terms of both probability estimation and community detection in almost all examples, except that its probability estimation accuracy is comparable to JIM in Example 1 with strong-balanced signed networks. Note that JIM is not applicable to detect communities in weak-balanced signed networks with $K\geq 2$, and thus the blanks in Tables 3 and 5. Its probability estimation accuracy also becomes substantially less competitive than SBBM in Examples 2 and 3 with weak-balanced signed networks. Furthermore, across all examples, the community detection errors and probability estimation errors of SBBM decrease significantly with $n$, confirming its asymptotic consistencies in Theorems 2 and 3. In contrast, the performance of SLP and SPONGE improves slowly with $n$ in Examples 2 and 3 with node heterogeneity, which is possibly due to the fact that node heterogeneity is largely ignored by SLP and SPONGE. Particularly, in Example 3, the performance of SBBM becomes better as $\sigma$ increases or node heterogeneity becomes stronger, yet such a positive trend is not observed in the performances of SLP and SPONGE. This again confirms the capability of SBBM in better capturing node heterogeneity than the other methods.
5.2 International relation network
We now turn to apply SBBM to analyze an international relation network constructed from two sources, the BACI trade database (CEPII, version 2025-01) [CEPII:2010-23], which harmonizes annual bilateral trade flows, and the Global Sanctions Data Base R4 [felbermayr2020global; yalcin2025global] that catalogs economic sanctions between economies. For each pair of economies, if economic sanctions have ever existed between them during 2015-2023, a negative edge is assigned; otherwise, a positive edge is assigned if their bilateral trade volume ranks in the top $1/30$. This pre-processing leads to a signed network with 164 economies, 580 positive edges, and 589 negative edges. It contains 2,756 strong-balanced triads, 3,559 weak-balanced triads, and 847 unbalanced triads. It is evident that weak-balanced triads occur more than four times as frequently as unbalanced ones, in sharp contrast to the expected ratio 5:3 under random assignment. This substantial deviation from randomness indicates a pronounced weak balance structure in the international relation network.
The community structure in the international relation network is then estimated via SBBM, SLP, SPONGE and JIM, respectively. Also, following the suggestion in zhang2022signed, the number of communities $K=4$ is determined by optimizing the Bayesian information criterion [saldana2017many]. For comparison, we compute the signed modularity gomez2009analysis of estimated communities from different methods. Particularly, for each $i\in[n]$, let $k^{+}_{i}=\sum_{j\in[n]}a^{+}_{ij}$, $k^{-}_{i}=\sum_{j\in[n]}a^{-}_{ij}$, $m^{+}=\frac{1}{2}\sum_{i,j\in[n]}a^{+}_{ij}$, and $m^{-}=\frac{1}{2}\sum_{i,j\in[n]}a^{-}_{ij}$. The signed modularity of a given community assignment $\{\psi_{i}\}_{i\in[n]}$ is defined as
$Q_{\text{signed}}=\frac{1}{2m^{+}}\sum_{i,j\in[n]}\left(a^{+}_{ij}-\frac{k_{i}^{+}k_{j}^{+}}{2m^{+}}\right)\delta(\psi_{i},\psi_{j})-\frac{1}{2m^{-}}\sum_{i,j\in[n]}\left(a^{-}_{ij}-\frac{k_{i}^{-}k_{j}^{-}}{2m^{-}}\right)\delta(\psi_{i},\psi_{j}),$
where $\delta(\psi_{i},\psi_{j})=1$ if $\psi_{i}=\psi_{j}$ and 0 otherwise. Note that $Q_{\text{signed}}$ takes values in [2,2][-2,2], and a larger value indicates more positive edges within communities and more negative edges between communities. Table 7 reports the signed modularities of all methods, showing that SBBM achieves the largest value, thereby demonstrating its advantages against the other methods.
Table 7: Signed modularity of various methods on the international relation network.
SBBM SLP SPONGE JIM
$Q_{\text{signed}}$ 0.4266 0.1943 0.1296 -0.0032
Figure 2 further displays the heatmap of the international relation network, where nodes are reordered according to $\widehat{\bm{Z}}$ from SBBM, and positive and negative edges are dyed into red or blue. It is clear that most positive edges occur within the same community, while negative edges occur across different communities, and a particularly high concentration of positive edges can be observed in the largest community.
Figure 2: Heatmap of the international relation network.
We also plot the estimated communities of SBBM on a world map in Figure 3. The sizes of the four estimated communities are 106, 29, 24, and 5, respectively. The largest one is centered on China and mainly consists of developing economies in Asia, Africa and Latin America. The second community is centered on the United States and mainly consists of developed economies in Europe and North America. The third community mainly consists of Afghanistan, Belarus, Iran, Syria, North Korea, Russia, Sudan and Lebanon, and the fourth community consists of Bosnia and Herzegovina, Cyprus, Eritrea, Moldova, and Ukraine. The economies in these two communities are conflict-affected or frequently subject to sanctions from other economies, and more interestingly, there are even stronger sanction ties between them. We further scrutinize the patterns of positive and negative edges via the estimated parameters. It is noteworthy that both China and the United States have high $\widehat{\beta}^{+}$, indicating their leading roles in forming strong within-community trading relations. On the other hand, both the United States and the United Kingdom have high $\widehat{\gamma}^{-}$, indicating their leading roles in imposing sanctions.
Figure 3: The four estimated communities by SBBM on the world map, where economics in different communities are dyed with different colors.
6 Summary
This paper proposes a novel signed block $\beta$-model to model the strong and weak balance structures in signed networks simultaneously. It introduces four sets of parameters that distinguish each node’s connection tendencies within its community and across different communities, which greatly facilitates community detection in signed networks with balance structures and node heterogeneity. We establish the identifiability conditions for the signed block $\beta$-model, and develop an efficient alternative updating algorithm for parameter estimation and community detection. Its asymptotic properties are also established in terms of both probability estimation and community detection, which are further supported by numerical experiments on both synthetic and real-world experiments. Note that statistical inference about the signed block $\beta$-model is also of interest but possibly hindered by the discrete feature of the community structure, and we leave it to future investigation.
Acknowledgment
JZ’s research is supported in part by National Key R&\&D Program of China (2024YFA1012200), “USTC Research Funds of the Double First-Class Initiative” (YD2040002020) and NSFC-12301388, and JW’s research is supported in part by HK RGC Grants GRF-11311022, GRF-14306523, GRF-14303424, and CUHK Startup Grant 4937091.

### v2

@@
 In this paper, we formally define strong and weak balance in signed networks, and propose a signed block $\beta$-model, which is capable of modeling strong- and weak-balanced signed networks simultaneously.
-We establish the identifiability of the proposed model by leveraging properties of bipartite graphs, and develop an efficient alternating updating algorithm to optimize the resulting log-likelihood function.
-More importantly, we establish the asymptotic consistencies of the proposed model in terms of both probability estimation and community detection.
+We establish the identifiability of the proposed model by leveraging properties of bipartite graphs, and develop an efficient two-step algorithm to tackle the resulting optimization problem and recover the community membership.
+More importantly, we establish the asymptotic consistencies of the proposed model in terms of both parameter estimation and community detection.
 Its advantages are also demonstrated through extensive numerical experiments and the application to a real-world international relationship network.
@@
 1 Introduction
-Network data has attracted increasing attention in recent years, with applications in diverse scientific domains, ranging from social science [hunter2008goodness], computer science [heard2014filtering] to biomedicine [shojaie2009analysis].
-Analysis of network data can reveal latent structures underlying the interactions among entities of interest, such as community detection [girvan2002community], link prediction [martinez2016survey], homophily [hoff2007modeling] and so on.
+Network data has attracted increasing attention in recent years, with applications in diverse scientific domains, ranging from social science [Hunter et al., 2008], computer science [Heard et al., 2014] to biomedicine [Shojaie and Michailidis, 2009].
+Analysis of network data can reveal latent structures underlying the interactions among entities of interest, such as community detection [Girvan and Newman, 2002], link prediction [Martínez et al., 2016], homophily [Hoff, 2007] and so on.
 Most existing methods for network analysis focus on unsigned networks, whereas signed networks are also prevalent in real-world applications.
-Specifically, signed networks have been extensively employed to represent relationships among objects [harary1953notion], which associate polarity information with each edge and thus distinguish friends from enemies, or allies from rivals.
-These signed edges give rise to unique structural characteristics grounded in social psychology theories, particularly the balance theory [heider1946attitudes; cartwright1956structural], stating that certain triadic configurations involving signed edges lead to stable or unstable social structures.
+Specifically, signed networks have been extensively employed to represent relationships among objects [Harary, 1953], which associate polarity information with each edge and thus distinguish friends from enemies, or allies from rivals.
+These signed edges give rise to unique structural characteristics grounded in social psychology theories, particularly the balance theory [Heider, 1946; Cartwright and Harary, 1956], stating that certain triadic configurations involving signed edges lead to stable or unstable social structures.
 Figure 1 illustrates four possible triads, among which triads A and C represent strong-balanced structures.
 These triads align with intuitive social principles, such as “a friend of my friend is likely my friend” and “an enemy of my friend is likely my enemy”.
-Weak balance [davis1967clustering] extends this concept by additionally permitting triad D, which accommodates the occurrence of three mutually hostile nodes.
+Weak balance [Davis, 1967] extends this concept by additionally permitting triad D, which accommodates the occurrence of three mutually hostile nodes and is thus more suitable for modeling real-life signed networks.
 Figure 1: Configurations of four possible triads.
-Community detection has been widely studied for unsigned networks [holland1983stochastic; girvan2002community; newman2006modularity; lei2015consistency; zhang2022directed], but much less so for signed networks.
-Most existing community detection methods for signed networks [doreian1996partitioning; kunegis2010spectral; jiang2015stochastic; cucuringu2021regularized] are algorithm based and lack formal theoretical analysis, not to mention that the prevalent balanced structures in signed networks are often ignored by the aforementioned community detection methods.
-Note that the balance theory actually posits that a strong-balanced (weak-balanced) network can be partitioned into two (multiple) sub-networks such that positive edges connect nodes within the same sub-network, whereas nodes from different sub-networks are connected by negative edges [easley2010networks].
-More recently, tang2024population introduces a formal definition of strong-balanced signed networks in a probabilistic sense and proposes an efficient method to detect the two communities, yet the method is not readily applicable to the weak-balanced signed networks with multiple communities.
-In this paper, we first extend the definition of strong-balanced signed networks in tang2024population to accommodate weak-balanced signed networks.
-We also provide a set of equivalence conditions for both strong- and weak-balanced signed networks, from both local and global perspectives.
+Community detection has been widely studied for unsigned networks [Holland et al., 1983; Girvan and Newman, 2002; Newman, 2006; Lei and Rinaldo, 2015; Zhang et al., 2022a], but much less so for signed networks.
+Most existing community detection methods for signed networks [Kunegis et al., 2010; Jiang, 2015; Cucuringu et al., 2021] are algorithm based and lack formal theoretical analysis, not to mention that the prevalent balanced structures in signed networks are often ignored by the aforementioned community detection methods.
+Note that the balance theory actually posits that a strong-balanced (weak-balanced) network can be partitioned into two (multiple) sub-networks such that positive edges connect nodes within the same sub-network, whereas nodes from different sub-networks are connected by negative edges [Easley et al., 2010].
+More recently, tang2024population introduces a formal definition of strong-balanced signed networks in a probabilistic sense and proposes an efficient method to detect the two communities, yet the method is inapplicable to analyze the more realistic weak-balanced signed networks with multiple communities.
+In this paper, we first provide a new set of conditions for both strong- and weak-balanced signed networks, from both local and global perspectives.
 Built upon the global perspective, we then propose a signed block $\beta$-model to explicitly incorporate the strong or weak balance structures in heterogeneous signed networks.
 Particularly, it introduces two sets of parameters to capture each node’s distinct connection tendencies within its own community and across different communities.
-The identifiability of the signed block $\beta$-model is established by exploiting properties of the bipartite graph, and an efficient alternating updating algorithm is developed to optimize the resulting negative log-likelihood function for parameter estimation.
-More importantly, the asymptotic consistencies of the block $\beta$-model are established in terms of both probability estimation and community detection.
+The identifiability of the signed block $\beta$-model is established by exploiting properties of the bipartite graph, and an efficient two-step algorithm is developed to solve the resulting optimization problem and recover the community membership.
+More importantly, the asymptotic consistencies of the signed block $\beta$-model are established in terms of both parameter estimation and community detection.
 Its advantages are also supported in numerical experiments on various synthetic signed networks and a real-world international relationship network.
@@
 Such a treatment not only incorporates community structure in each node’s connection tendency, but also admits more substantial node heterogeneity than existing models [chen2021analysis; shao20212; cucuringu2021regularized].
-Moreover, a novel set of identifiability conditions for the proposed signed block $\beta$-model is established by leveraging techniques from the graph theory.
+In addition, a novel set of identifiability conditions for the proposed signed block $\beta$-model is established by leveraging techniques from the graph theory.
 The theoretical techniques developed here can be applicable to identifiability analyses of other network models involving community structures, and thus are of independent interest.
+Moreover, we incorporate the $(1+\epsilon)$-approximation algorithm from the literature of projective clustering har2002projective; deshpande2006matrix; tukan2022new into the second step of our method to solve the $K$-Lines clustering problem, namely, to compute $K$ lines to approximate a set of points.
+Although li2024dissecting discusses the $K$-Lines clustering problem, their algorithm lacks theoretical guarantee.
 The remainder of this paper is organized as follows.
 Section 2 formally gives the definition of strong and weak balance in signed networks.
-Section 3 presents the signed block $\beta$-model, establishes its identifiability conditions, and develops an efficient alternative updating scheme to tackle the resulting optimization task.
-Section 4 provides a rigorous theoretical analysis of the proposed signed block $\beta$-model and quantifies its asymptotic behaviors in both probability estimation and community detection.
+Section 3 presents the signed block $\beta$-model, establishes its identifiability conditions, and develops an efficient two-step algorithm to tackle the resulting optimization task.
+Section 4 provides a rigorous theoretical analysis of the proposed signed block $\beta$-model and quantifies its asymptotic behaviors in both parameter estimation and community detection.
 Section 5 examines the numeric performance of the signed block $\beta$-model via extensive experiments on both synthetic and real-world signed networks.
 Finally, Section 6 concludes the paper with a brief summary.
-Technical proofs of the main theorems are provided in the Appendix, whereas auxiliary lemmas are contained in the Supplementary Material.
+Technical proofs of the main theorems and propositions are provided in the Appendix, whereas auxiliary lemmas are contained in the Supplementary Material.
 Notations.
@@
 For any vector $\bm{v}=(v_{1},\dots,v_{n})^{\top}\in\mathcal{R}^{n}$, $\|\bm{v}\|_{2}^{2}=\sum_{i\in[n]}v_{i}^{2}$, and $\|\bm{v}\|_{\infty}=\max_{i\in[n]}|v_{i}|$.
-We use $\bm{1}_{n}$ to denote the $n$-dimensional column vector with all elements being 1 and $\mathbb{I}(\cdot)$ to denote the indicator function.
-For any matrix $\bm{A}=(a_{ij})\in\mathcal{R}^{n\times m}$, let $\bm{A}_{i}$ be its $i$-th row, and $\|\bm{A}\|_{F}^{2}=\sum_{i\in[n],j\in[m]}a_{ij}^{2}$ be its Frobenius norm.
+We use $\bm{1}_{n}$ to denote the $n$-dimensional column vector with all elements being 1 and $\mathbf{1}(\cdot)$ to denote the indicator function.
+For any matrix $\bm{A}=(a_{ij})\in\mathcal{R}^{n\times m}$, let $\bm{A}_{i\cdot}$ be its $i$-th row, $\bm{A}_{\cdot i}$ be its $i$-th column, and $\|\bm{A}\|_{F}^{2}=\sum_{i\in[n],j\in[m]}a_{ij}^{2}$ be its Frobenius norm.
 For any two matrices $\bm{A}$ and $\bm{B}$, $\bm{A}\odot\bm{B}$ is the Hadamard product with $(\bm{A}\odot\bm{B})_{ij}=a_{ij}b_{ij}$.
+Let $\bm{J}=I_{n}-\frac{\bm{1}_{n}\bm{1}_{n}^{\top}}{n}$ be the centering matrix.
+For any two vectors $\bm{x}=(x_{1},\dots,x_{n})$ and $\bm{y}=(y_{1},\dots,y_{n})$, $\bm{x}<\bm{y}$ denotes $x_{i}<y_{i}$ for each $i\in[n]$.
 For any two sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we denote $a_{n}\prec b_{n}$ if $a_{n}=o(b_{n})$, $a_{n}\preceq b_{n}$ if $a_{n}=O(b_{n})$ and $a_{n}\simeq b_{n}$ if $a_{n}\preceq b_{n}$ and $b_{n}\preceq a_{n}$.
@@
 For any two sets $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{X}\setminus\mathcal{Y}=\{x\mid x\in\mathcal{X}\text{ and }x\notin\mathcal{Y}\}$.
-2 Signed network and balance theory
+2 Signed network and balance structure
 A signed network $\mathcal{G}$ can be represented via its symmetric adjacency matrix $\bm{A}=(a_{ij})_{i,j\in[n]}$ with $a_{ij}\in\{1,-1,0\}$, where $a_{ij}=1$ indicates a positive edge between nodes $i$ and $j$, $a_{ij}=-1$ indicates a negative edge and $a_{ij}=0$ indicates no edge at all.
@@
 One of the most salient features of signed network is the widely observed balance structure [harary1953notion; davis1967clustering], positing that relationships among entities tend toward balanced sign configurations.
-Definition 1 (Strong balance)
+Definition 1 (Strong balance).
 A signed network is strong-balanced if for any triple $(i,j,k)$,
@@
 Definition 1 assures that a strong-balanced signed network does not contain a triad achieving the highest probability at triad B or D as shown in Figure 1.
-An equivalent definition has also been introduced in tang2024population, which quantifies the probabilistic strong balance for signed networks.
-Moreover, Definition 1 can be extended to define weak-balanced signed networks.
-Definition 2 (Weak balance)
+It can also be extended to define weak-balanced signed networks.
+Definition 2 (Weak balance).
 A signed network is weak-balanced if for any triple $(i,j,k)$,
@@
 It only restricts the occurrence of triad B, and thus aligns with the non-probabilistic definitions of weak balance in the existing literature [easley2010networks].
-Proposition 1
+Proposition 1.
 A signed network $\mathcal{G}$ is strong-balanced if and only if
@@
 (ii) There exists a partition $n]=C_{1}\cup C_{2}$ such that for each $i,j\in[n]$, $p_{ij}^{+}>p_{ij}^{-}$ if $i,j$ belong to the same subset, and $p_{ij}^{+}<p_{ij}^{-}$, if $i,j$ belong to different subsets.
-Proposition 2
+Proposition 2.
 A signed network $\mathcal{G}$ is weak-balanced if and only if
@@
 Propositions 1 and 2 characterize the strong and weak balance structures in signed networks from both local and global perspectives.
-In fact, the local perspective (i) in Proposition 1 is essentially equivalent to $\mathbb{E}(a_{ij}a_{jk}a_{ki}\big||a_{ij}a_{jk}a_{ki}|=1)>0$ for any $(i,j,k)$, which is exactly the definition of the strong-balanced signed network in tang2024population, whereas (i) in Proposition 2 can be regarded as its analogy for the weak-balanced signed network.
-On the other hand, (ii) in Proposition 1 and (ii) in Proposition 2 give the global perspectives on strong and weak balance structures, which greatly facilitate the modeling of community structures in signed networks.
+In fact, the local perspective (i) in Proposition 1 is essentially equivalent to $\mathbb{E}(a_{ij}a_{jk}a_{ki}\big||a_{ij}a_{jk}a_{ki}|=1)>0$ for any $(i,j,k)$, which is exactly the definition of the strong-balanced signed network in tang2024population, whereas Proposition 2(i) can be regarded as its analogy for the weak-balanced signed network.
+On the other hand, Proposition 1(ii) and 2(ii) give the brand new global perspectives on strong and weak balance structures based on partitions of the signed network, which greatly facilitate the modeling of community structures in signed networks.
 3 Signed block $\beta$-model
 We propose a signed block $\beta$-model, denoted as SBBM, to leverage the global perspectives of the balance structures in signed networks.
-Particularly, let $\psi_{i}\in[K]$ denote the community membership of node $i$, then
+Particularly, let $\sigma:[n]\rightarrow[K]$ denote the community membership of node $i$, then
 […]
@@
 It follows from Propositions 1 and 2 that SBBM with constraints (2) satisfies strong balance when $K=2$ and weak balance when $K\geq 3$.
+In contrast, enforcement of strong and weak balance structures can be challenging in most other existing signed network models.
 3.1 Model identifiability
-Let $\bm{Z}=(\bm{z}_{1},\dots,\bm{z}_{n})^{\top}$ be the community membership matrix, where $\bm{z}_{i}\in\{0,1\}^{K}$ with $z_{ik}=1$ if $\psi_{i}=k$.
+Let $\bm{Z}=(\bm{z}_{1},\dots,\bm{z}_{n})^{\top}$ be the community membership matrix, where $\bm{z}_{i}\in\{0,1\}^{K}$ with $z_{ik}=1$ if $\sigma(i)=k$.
 Define $\bm{\beta}^{+}=(\beta^{+}_{i})_{i=1}^{n}$, $\bm{\gamma}^{+}=(\gamma_{i}^{+})_{i=1}^{n}$, $\bm{\beta}^{-}=(\beta^{-}_{i})_{i=1}^{n}$, $\bm{\gamma^{-}}=(\gamma_{i}^{-})_{i=1}^{n}$, $\bm{\eta}^{+}=\bm{\beta}^{+}-\bm{\gamma}^{+}$ and $\bm{\eta}^{-}=\bm{\beta}^{-}-\bm{\gamma}^{-}$.
 Then, the parameter space is
-$\mathcal{P}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ $\times\mathbb{R}^{n}\times\{0,1\}^{n\times n}:$
+$\mathcal{P}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ $\times\mathbb{R}^{n}\times\{0,1\}^{n\times K}:$
 […]
@@
 We say SBBM is identifiable if parameters $(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}\bm{Z}^{\top})$ can be uniquely determined from $\bm{\Theta}^{+}$ and $\bm{\Theta}^{-}$, where $\bm{Z}$ can be uniquely determined up to column permutation.
-Theorem 1
+Theorem 1.
 SBBM is identifiable when $K=2$, if $\eta_{1}^{+}=0$, $\eta_{1}^{-}=0$ and there does not exist $S_{1}\subseteq C_{1}$ and $S_{2}\subseteq C_{2}$ such that $\eta_{i}^{+}+\eta_{j}^{+}=0$ or $\eta_{i}^{-}+\eta_{j}^{-}=0$ for any $i\in S_{1},j\in S_{2}$ and $i\in C_{1}\setminus S_{1},j\in C_{2}\setminus S_{2}$.
@@
 3.2 Computing algorithm
-The negative log-likelihood function of SBBM is
-$L(\bm{\Theta};{\bm{A}})=-\frac{2}{n(n+1)}\sum_{1\leq i\leq j\leq n}\Big(a^{+}_{ij}\theta^{+}_{ij}+a^{-}_{ij}\theta^{-}_{ij}-\log\big(1+e^{\theta^{+}_{ij}}+e^{\theta^{-}_{ij}}\big)\Big),$ (3)
-where $\bm{\Theta}=(\bm{\Theta}^{+},\bm{\Theta}^{-})$, $a_{ij}^{+}=\mathbb{I}(a_{ij}=1)$, $a_{ij}^{-}=\mathbb{I}(a_{ij}=-1)$ for each $1\leq i\leq j\leq n$. Define
-$\mathcal{P}_{0}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathcal{P}:\$$\eta_{i}^{+}+\eta_{j}^{+}=0\ \text{or}\ \eta_{i}^{-}+\eta_{j}^{-}=0$
-$\text{for any}\ i\in S_{1},j\in S_{2}\ \text{and}\ i\in S_{1}^{c},j\in S_{2}^{c}\big\},$
-where $S_{1}\subseteq C_{1}$ and $S_{2}\subseteq C_{2}$, and
-[…]
-Then, $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ can be obtained by minimizing $L(\bm{\Theta};{\bm{A}})$ over $(\mathcal{P}\setminus\mathcal{P}_{0})\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$ and over $\mathcal{P}\setminus\mathcal{P}_{1}$ when $K\geq 3$.
-It is clear that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ are zero sets under the Lebesgue measure.
-Therefore, we get $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ by minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$, and minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}$ directly when $K\geq 3$ [latouche2011overlapping; zhang2022identifiability].
-We develop an alternative updating scheme to optimize $L(\bm{\Theta};{\bm{A}})$.
-We first rewrite $L(\bm{\Theta};\bm{A})$ as $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})$ and initialize $\bm{\gamma}^{+(0)}=\bm{\eta}^{+(0)}=\bm{\gamma}^{-(0)}=\bm{\eta}^{-(0)}=\bm{0}_{n}$ and $\bm{Z}^{(0)}$ by applying spectral clustering algorithm to $\bm{A}$.
-Given $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})$ at the tt‑th step for $t\geq 1$, note that, with $\bm{Z}^{(t-1)}$ fixed, minimizing $L(\bm{\Theta};\bm{A})$ subject to (2) with respect to $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)})$ is a convex optimization problem.
-We therefore adopt the Spectral Projected Gradient method [birgin2000nonmonotone], which combines a Barzilai–Borwein spectral step, Euclidean projection onto the feasible set, and an Armijo back-tracking line search, thereby handling curvature adaptation and step-size control automatically.
-Upon convergence, we obtain the updated parameters $(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)})$.
-Note that we project $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ such that $\eta_{1}^{+(t)}=\eta_{1}^{-(t)}=0$ when $K=2$, which are still denoted as $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ with a slight abuse of notation.
-Next, we adopt the similar method as in yuan2021community and zhang2024consistent to update $\bm{Z}^{(t)}$ due to its discrete nature.
-Let $\bm{Z}_{-i}$ be a matrix by deleting of the ii‑th row of $\bm{Z}$.
-Then, node $i$ will be more likely to belong to the kk‑th community if $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{ik}=1)\leq L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{il}=1)$ for each $l\neq k$.
-Thus, we have
-$z_{ik}^{(t)}=\begin{cases}1,&k=\displaystyle\arg\min_{l\in[K]}L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}_{-i}^{(t-1)},z_{il}=1\right);\\ 0,&\text{otherwise},\end{cases}$
-for each $i\in[n]$ and $k\in[K]$.
-After updating $\bm{Z}^{(t)}$, we check whether
-$L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)}\right)<L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t-1)}\right).$
-Otherwise, we update each $\bm{z_{i}}$ sequentially, thereby ensuring that the negative log‑likelihood always decreases.
-Let $\alpha$ and $T_{\max}$ be threshold values for the relative error and updating steps.
-The whole algorithm will stop when $t>T_{\max}$ or the relative error
-$\frac{\bigl|L(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)})-L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}{\bigl|L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}<\alpha.$
-Let $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ be the output of the alternative updating scheme, then $\widehat{\bm{\Theta}}^{+}=\widehat{\bm{\gamma}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{+\top}+(\widehat{\bm{\eta}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{+\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$ and $\widehat{\bm{\Theta}}^{-}=\widehat{\bm{\gamma}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{-\top}+(\widehat{\bm{\eta}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{-\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$.
+The negative log-likelihood of $\bm{A}^{+}=(a^{+}_{ij})_{i,j=1}^{n}$ and $\bm{A}^{-}=(a^{-}_{ij})_{i,j=1}^{n}$ can be written as
+$\mathcal{L}_{n}(\bm{\Theta}^{+},\bm{\Theta}^{-})=-\sum_{1\leq i,j\leq n}\Big(a^{+}_{ij}\theta^{+}_{ij}+a^{-}_{ij}\theta^{-}_{ij}-\log\big(1+e^{\theta^{+}_{ij}}+e^{\theta^{-}_{ij}}\big)\Big),$
+where $a_{ij}^{+}=\mathbb{I}(a_{ij}=1)$ and $a_{ij}^{-}=\mathbb{I}(a_{ij}=-1)$. Then we have
+$(\widehat{\bm{\Theta}}^{+},\widehat{\bm{\Theta}}^{-})=\operatorname*{arg\,min}_{\bm{\Theta}^{+},\bm{\Theta}^{-}}\left\{\mathcal{L}_{n}(\bm{\Theta}^{+},\bm{\Theta}^{-})+\lambda_{n}^{+}\|\bm{\Theta}^{+}\|_{*}+\lambda_{n}^{-}\|\bm{\Theta}^{-}\|_{*}\right\},$ (3)
+where $\|\cdot\|_{*}$ denotes the nuclear norm, which is used to promote the low-rank structure in $\bm{\Theta}^{+}$ and $\bm{\Theta}^{-}$, and $\lambda_{n}^{+}$ and $\lambda_{n}^{-}$ are two tuning parameters.
+The objective function in (3) is convex, and it can be efficiently optimized via the accelerated proximal gradient algorithm [beck2009fast].
+The algorithm iterates between a gradient descent step on the smooth loss $\mathcal{L}_{n}$ and a proximal step for the nuclear norm, which is computed via the singular value thresholding operator [cai2010singular], $\mathcal{D}_{\tau}(\bm{M})=\sum_{k}\max\{\sigma_{k}-\tau,0\}\bm{u}_{k}\bm{v}_{k}^{\top}$ for any matrix $\bm{M}$ with singular value decomposition $\bm{M}=\sum_{k}\sigma_{k}\bm{u}_{k}\bm{v}_{k}^{\top}$.
+Once $\widehat{\bm{\Theta}}^{+}$ and $\widehat{\bm{\Theta}}^{-}$ are obtained, we then proceed to estimate $\sigma$.
+Proposition 3.
+Let $\bm{\Theta}_{\Delta}=\bm{\Theta}^{+}-\bm{\Theta}^{-}$ and $\bm{J}\bm{\Theta}_{\Delta}\bm{J}=\bm{Y}\bm{\Sigma}\bm{Y}^{\top}$ be the eigen decomposition with $\bm{Y}\in\mathbb{R}^{n\times r}$ and $r=\text{rank}\ (\bm{J}\bm{\Theta}_{\Delta}\bm{J})$.
+Then there are $K$ lines in $\mathbb{R}^{r}$, denoted as $\{l_{k}:\bm{w}_{k}+t\bm{v}_{k},t\in\mathbb{R},\bm{w}_{k},\bm{v}_{k}\in\mathbb{R}^{r},k=1,\dots,K\}$, such that each row of $\bm{Y}$ lies exactly on one of these $K$ lines.
+Furthermore, for any $i,j\in[n]$, if $\sigma(i)=\sigma(j)\in[K]$, then $\bm{Y}_{i\cdot}$ and $\bm{Y}_{j\cdot}$ live on the same line.
+Proposition 3 implies that the community structure in $\sigma$ can be estimated from the spectral decomposition of $\bm{J}\widehat{\bm{\Theta}}_{\Delta}\bm{J}$, where $\widehat{\bm{\Theta}}_{\Delta}=\widehat{\bm{\Theta}}^{+}-\widehat{\bm{\Theta}}^{-}$.
+Denote the eigenvectors of $\bm{J}\widehat{\bm{\Theta}}_{\Delta}\bm{J}$ as $\widehat{\bm{Y}}=(\widehat{\bm{Y}}_{1},\ldots,\widehat{\bm{Y}}_{n})^{\top}$, then we can recover $\sigma$ via a $K$-lines clustering of $\widehat{\bm{Y}}$.
+Formally, given $\{\widehat{\bm{Y}}_{1},\ldots,\widehat{\bm{Y}}_{n}\}$ and a set of lines $\bm{L}=\{\ell_{1},\dots,\ell_{K}\}$, we define
+$\Phi(\widehat{\bm{Y}},\bm{L},\sigma)=\frac{1}{n}\sum_{i=1}^{n}\|\widehat{\bm{Y}}_{i}-P_{\ell_{\sigma(i)}}\widehat{\bm{Y}}_{i}\|_{2}^{2},$
+where $P_{l}$ indicates the projection onto the $l$-th line.
+Let $\sigma_{\bm{L}}(i)=\arg\min_{k\in[K]}\|\widehat{\bm{Y}}_{i}-P_{l_{k}}\widehat{\bm{Y}}_{i}\|_{2}$, then $\Phi(\bm{Y},\bm{L},\sigma_{\bm{L}})=\min_{\sigma}\Phi(\bm{Y},\bm{L},\sigma)$.
+Therefore, the $K$-line clustering aims to find $\widehat{\bm{L}}=\arg\min_{\bm{L}}\Phi(\widehat{\bm{Y}},\bm{L},\sigma_{\bm{L}})$, which becomes a projective clustering problem [tukan2022new].
+The general projective clustering problem is NP-hard, but various techniques have been developed to obtain its $(1+\epsilon)$-approximated solutions [feldman2006coresets; ding2012linear; feldman2020turning].
+In particular, we adopt the method in ding2012linear, which guarantees an objective value no greater than $(1+\epsilon)$ times the optimal value over the entire search space.
+Denote the solution as $\widehat{\bm{L}}=\{\widehat{\ell}_{1},\dots,\widehat{\ell}_{K}\}$ such that
+$\Phi(\widehat{\bm{Y}},\widehat{\bm{L}},\sigma_{\widehat{\bm{L}}})\leq(1+\epsilon)\,\min_{\bm{L}}\Phi(\widehat{\bm{Y}},\bm{L},\sigma_{\bm{L}}),$
+and then the community membership is estimated as $\widehat{\sigma}=\sigma_{\widehat{\bm{L}}}$.
 4 Theory
-In this section, we establish the asymptotic properties of the proposed SBBM in terms of both probability estimation and community detection.
-Denote the true parameters as $(\bm{\beta}^{*+},\bm{\beta}^{*-},\bm{\gamma}^{*+},\bm{\gamma}^{*-},\bm{Z}^{*})\in\mathcal{P}$ with $\bm{\beta}^{*+}>\bm{\beta}^{*-}$ and $\bm{\gamma}^{*+}<\bm{\gamma}^{*-}$, and $\bm{\Theta}^{*}=(\bm{\Theta}^{*+},\bm{\Theta}^{*-})$ correspondingly.
-4.1 Consistency in probability estimation
-For each $i,j\in[n]$, denote the negative log-likelihood of $a_{ij}$ as
-$L(\theta_{ij};a_{ij})=-(a_{ij}^{+}\theta_{ij}^{+}+a_{ij}^{-}\theta_{ij}^{-})+\log\big(1+e^{\theta_{ij}^{+}}+e^{\theta_{ij}^{-}}\big),$
-where $\theta_{ij}=(\theta_{ij}^{+},\theta_{ij}^{-})$, and $L(\bm{\Theta};\bm{A})=\frac{2}{n(n+1)}\sum_{i\leq j}L(\theta_{ij};a_{ij})$.
-Let $e_{L}(\bm{\Theta},\bm{\Theta}^{*})=\mathbb{E}\big(L(\bm{\Theta};\bm{A})-L(\bm{\Theta}^{*};\bm{A})\big)$, which serves as an intermediate error metric measuring the accuracy of $\bm{\Theta}$ in estimating $\bm{\Theta}^{*}$.
-Assumption 1
-There exist constants $M_{1}<M_{2}$ such that both $\theta_{ij}^{+}$ and $\theta_{ij}^{-}$ are contained in $M_{1},M_{2}]$ for each $i,j\in[n]$.
+In this section, we establish the asymptotic properties of the proposed SBBM in terms of the estimation error of $(\widehat{\bm{\Theta}}^{+},\widehat{\bm{\Theta}}^{-})$ obtained in (3) and the community assignment error of $\widehat{\sigma}$ obtained in Section 3.2.
+4.1 Consistency in parameter estimation
+We first introduce several auxiliary quantities to facilitate the theoretical analysis.
+The true parameters are denoted by $\bm{\beta}^{*+},\bm{\beta}^{*-},\bm{\gamma}^{*+},\bm{\gamma}^{*-}$ and $\sigma^{*}$ with $\bm{\beta}^{*+}>\bm{\beta}^{*-}$ and $\bm{\gamma}^{*+}<\bm{\gamma}^{*-}$ component-wisely.
+Define $\bm{\Theta}^{*}=(\bm{\Theta}^{*+},\bm{\Theta}^{*-})$ and let $\bm{\Theta}_{\Delta}^{*}=\bm{\Theta}^{*+}-\bm{\Theta}^{*-}$.
+Finally, let $p^{*+}_{ij}=e^{\theta_{ij}^{*+}}/(1+e^{\theta_{ij}^{*+}}+e^{\theta_{ij}^{*-}})$, $p^{*-}_{ij}=e^{\theta_{ij}^{*-}}/(1+e^{\theta_{ij}^{*+}}+e^{\theta_{ij}^{*-}})$ be the true edge probabilities and $\bm{P}^{*+}=(p^{*+}_{ij})_{i,j=1}^{n},\bm{P}^{*-}=(p^{*-}_{ij})_{i,j=1}^{n}$.
+Assumption 1.
+The parameter $\bm{\Theta}=(\bm{\Theta}^{+},\bm{\Theta}^{-})$ satisfies that $M_{1}\leq\bm{\Theta}^{+}_{ij},\bm{\Theta}^{-}_{ij}\leq-M_{2}$ for all $1\leq i,j\leq n$, where $M_{1}$ and $M_{2}$ are allowed to diverge with $n$.
 Assumption 1 is common in the literature of network modeling.
@@
 This is in sharp contrast to the homogeneous assumption on connecting probabilities in most existing literature on sparse networks, where $M_{1}$ and $M_{2}$ are essentially of the same magnitude [zhang2022directed; li2023statistical; lei2023bias; zhen2023community].
-Proposition 3
-For any $\bm{\Theta}$ that satisfies Assumption 1, it holds that
-$\frac{2}{n(n+1)}\sum_{i\leq j}\mathbb{E}[L(\theta_{ij},a_{ij})-L(\theta^{*}_{ij},a_{ij})]^{2}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}e_{L}(\bm{\Theta},\bm{\Theta}^{*}),$
-$\frac{1}{n^{2}}\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}e_{L}(\bm{\Theta},\bm{\Theta}^{*}).$
-Proposition 3 paves a bridge between the first and second moments of $L(\bm{\Theta},\bm{A})-L(\bm{\Theta}^{*},\bm{A})$, which is crucial for establishing the weak consistency of $\widehat{\bm{\Theta}}$ in terms of $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$.
-Furthermore, it also relates $\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ to $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$, leading to the strong consistency of $\widehat{\bm{\Theta}}$ established in the following Theorem 2.
-Theorem 2
-[…]
-$\ P\Big(\frac{1}{n^{2}}\big\|\widehat{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}\geq 48e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\Big)\leq 2\exp(-c_{1}Kn\log n),$
-when $n$ is large enough, provided that $\epsilon_{n}^{2}\simeq\frac{Ke^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}\log n}{n}\prec 1$.
-Theorem 2 guarantees that $\bm{\Theta}^{*}$ can be consistently estimated if $K\simeq 1$ and $\min\{2e^{M_{2}}+e^{2M_{2}},1\}e^{2\max\{|M_{1}|,|M_{2}|\}}n^{-1}\log n$ is vanishing.
-In particular, if $M_{2}-M_{1}\preceq 1$ and both of them diverge to -$\infty$, then it is required that $\min\{p_{ij}^{*+},p_{ij}^{*-}\}\succ\frac{\log n}{n}$ for each $i,j\in[n]$, which matches the best sparsity requirement in the literature [lei2015consistency; ma2020universal; zhang2022directed].
-More importantly, when $M_{2}-M_{1}\succ 1$, $p_{ij}^{*+}$ and $p_{ij}^{*-}$ can behave differently to accommodate edge heterogeneity.
-Specifically, letting $M_{1}\to$-$\infty$ and $M_{2}\to\infty$, Theorem 2 still holds when $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$, resulting in a heterogeneous case with $p_{ij}^{*+},p_{ij}^{*-}\in(n^{-1/2}(\log n)^{1/2},1)$ for each $i,j\in[n]$.
+Theorem 2.
+Suppose Assumption 1 holds and $\min\{\lambda_{n}^{+},\lambda_{n}^{-}\}\geq c_{0}\sqrt{\max\{ne^{-M_{2}},\log n\}}$ for a large enough constant $c_{0}$, it then holds true with probability at least $1-n^{-c}$ that
+$\frac{1}{n^{2}}\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}^{*}_{\Delta}\|^{2}_{F}\leq\frac{c^{\prime}K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{{\lambda_{n}^{+}}^{2},{\lambda_{n}^{-}}^{2}\}.$ (4)
+for some positive constants $c$ and $c^{\prime}$. Furthermore, if we set $\lambda_{n}^{+},\lambda_{n}^{-}\simeq\sqrt{\max\{ne^{-M_{2}},\log n\}}$, then with the same probability, we have
+$\frac{1}{n^{2}}\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}^{*}_{\Delta}\|^{2}_{F}\leq\frac{c^{\prime\prime}K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{ne^{-M_{2}},\log n\}.$ (5)
+for some positive constant $c^{\prime\prime}$.
+Theorem 2 guarantees that $\bm{\Theta}^{*}_{\Delta}$ can be consistently estimated if $\lambda^{+}$ and $\lambda^{-}$ are sufficiently large but $\frac{K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{{\lambda_{n}^{+}}^{2},{\lambda_{n}^{-}}^{2}\}$ vanishes.
+Its optimal convergence rate is attained when $\lambda_{n}^{+},\lambda_{n}^{-}\simeq\sqrt{\max\{ne^{-M_{2}},\log n\}}$.
+If $K\simeq 1$ and $M_{1}-M_{2}\preceq 1$ but both diverge to $\infty$, then it is required that $\min\{p_{ij}^{*+},p_{ij}^{*-}\}\succ\frac{\sqrt{\log n}}{n}$ for each $i,j\in[n]$, which matches the best sparsity requirement in the literature [lei2015consistency; ma2020universal; zhang2022directed].
+Furthermore, when $M_{1}-M_{2}\succ 1$, $p_{ij}^{*+}$ and $p_{ij}^{*-}$ can behave differently to accommodate edge heterogeneity.
+Specifically, letting $e^{-M_{1}}\prec 1$ and $\frac{\sqrt{\log n}}{n}\prec e^{-M_{2}}\preceq 1$, Theorem 2 still holds when $e^{-M_{1}}\succ n^{-1/2}$, resulting in a heterogeneous case with $p_{ij}^{*+},p_{ij}^{*-}\in(n^{-1/2},1)$ for each $i,j\in[n]$.
 This is consistent with the heterogeneous result for the latent space model [ma2020universal; zhang2022joint], and better than that for the classical $\beta$-model [shao20212], where the connecting probabilities can only vary in $(n^{-1/4},1)$.
 4.2 Consistency in community detection
-We next turn to establish the asymptotic properties of $\widehat{\bm{Z}}$, whose community detection error is defined as
-$d(\widehat{\bm{Z}},\bm{Z}^{*})=\min_{\bm{\Pi}}\frac{1}{n}\|\widehat{\bm{Z}}\bm{\Pi}-\bm{Z}^{*}\|_{F}^{2},$ (4)
-where minimization is taken over all permutation matrices $\bm{\Pi}\in\{0,1\}^{K\times K}$.
-The quantity $d(\widehat{\bm{Z}},\bm{Z}^{*})$ measures the minimal fraction of misclassified nodes, up to permutation, and is commonly used in the community detection literature [lei2015consistency].
-Assumption 2
-There exists a constant $M_{\gamma}$ such that $\|\bm{\gamma}\|_{\infty}\leq M_{\gamma}$.
-Let $\mathcal{S}_{i}^{*+}=\big\{j\in[n]:\ |\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}\big\}$ and $\mathcal{S}_{i}^{*-}=\big\{j\in[n]:\ |\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}\big\}$ for some $c_{\eta}>0$, then $\max_{i\in[n]}|\mathcal{S}_{i}^{*+}|\prec\frac{n}{K}$ or $\max_{i\in[n]}|\mathcal{S}_{i}^{*-}|\prec\frac{n}{K}$.
-Assumption 2 assures that the number of pairs of $(i,j)$ satisfying $|\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}$ or $|\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}$ is substantially less than $n/K$.
-Note that $|\beta_{i}^{*+}+\beta_{j}^{*+}|\geq 2M_{\gamma}+c_{\eta}$ and $|\beta_{i}^{*-}+\beta_{j}^{*-}|\geq 2M_{\gamma}+c_{\eta}$ imply that $\min\{|\eta_{i}^{*+}+\eta_{j}^{*+}|,|\eta_{i}^{*-}+\eta_{j}^{*-}|\}\geq c_{\eta}$, thus $c_{\eta}$ can be regarded as a measurement of the difference in the level of connection probability within the same community and those across different communities.
-For any $(\bm{\gamma}^{+},\bm{\gamma}^{-},\bm{\eta}^{+},,\bm{\eta}^{-},\bm{Z})\in\mathcal{P}$, let $\mathcal{N}_{k}=\{i\in[n]:z_{ik}=1\}$ and $n_{k}=|\mathcal{N}_{k}|$ for each $k\in[K]$.
-Proposition 4
-Suppose all assumptions in Theorem 2 and Assumption 2 hold, and that $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$.
-Then, with probability at least $1-2\exp(-c_{1}Kn\log n)$, there does not exist $k_{1}\neq k_{2}\in[K]$ and $k\in[K]$ such that $|\mathcal{N}_{k_{1}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq|\mathcal{N}_{k_{2}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq\frac{n}{K}$.
-The requirement of $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$ implies that $c_{\eta}$ can not be too small, which can be interpreted as a minimal signal strength condition.
-Proposition 4 assures that any estimated community $\widehat{\mathcal{N}}_{k}$ does not substantially overlap with more than one true community with high probability, which further leads to the estimation consistency of $\widehat{\bm{Z}}$ as shown in Theorem 3.
-Theorem 3
-Suppose all assumptions in Proposition 4 hold, then there exists a positive constant $c_{2}$ such that when $n$ is sufficiently large,
-$P\big(d(\widehat{\bm{Z}},\bm{Z}^{*})\geq c_{2}Kc_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\big)\leq 2\exp(-c_{1}Kn\log n).$
-Theorem 3 establishes the convergence rate of $d(\widehat{\bm{Z}},\bm{Z}^{*})$ for community detection.
-Particularly, in the homogeneous case with $\max\{|M_{1}|,|M_{2}|,K,c_{\eta}\}\simeq 1$, the convergence rate becomes $n^{-1}\log n$, which matches with the best community detection result in literature [lei2015consistency; zhang2022directed].
-In the heterogeneous case with $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$ where $p_{ij}^{*}$ is allowed to vary in $(n^{-1/2}(\log n)^{1/2},1)$ for each $i\leq j$, $c_{\eta}$ can diverge to infinity at a rate no faster than $\log n$ and Theorem 3 continues to hold with $d(\widehat{\bm{Z}},\bm{Z}^{*})=O_{p}\big((\log n)^{-2}\big)$.
+Let $K\simeq 1$ and $\frac{1}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{ne^{-M_{2}},\log n\}\prec 1$. We now turn to establish the asymptotic properties of $\widehat{\sigma}$. Let
+$d(\widehat{\sigma},\sigma^{*})=\min_{\pi}\frac{1}{n}\sum_{i\in[n]}\big(\pi\circ\widehat{\sigma}(i)-\sigma^{*}(i)\big)^{2}$ (6)
+be the community detection error, where $\pi:[K]\rightarrow[K]$ is the permutation mapping.
+Then, $d(\widehat{\sigma},\sigma^{*})$ measures the minimal fraction of misclassified nodes up to permutation, and is commonly used in the literature [lei2015consistency].
+Assumption 2.
+Let $\bm{\eta}_{\Delta}^{*}=(\bm{\beta}^{*+}-\bm{\gamma}^{*+})-(\bm{\beta}^{*-}-\bm{\gamma}^{*-})=(\eta_{\Delta,i}^{*})_{i=1}^{n}$.
+For each $k\in[K]$, let $s_{k}^{*}=\sum_{\sigma^{*}(i)=k}1$, $t_{k}^{*}=\sum_{\sigma^{*}(i)=k}\eta_{\Delta,i}^{*}$ and $u_{k}^{*}=\sum_{\sigma^{*}(i)=k}(\eta_{\Delta,i}^{*})^{2}$, then there exist $\bar{s}_{k}>0$ and $\bar{t}_{k},\bar{u}_{k}\in\mathbb{R}$ and $\epsilon_{n}\prec 1$ such that $\max_{k}\{|s_{k}^{*}/n-\bar{s}_{k}|,|t_{k}^{*}/n-\bar{t}_{k}|,|u_{k}^{*}/n-\bar{u}_{k}|\}\leq\epsilon_{n}$.
+Moreover, elements in $\{\bar{t}_{k}\pm\sqrt{\bar{s}_{k}\bar{u}_{k}},k\in[K]\}$ are all different.
+Assumption 2 imposes an empirical moments convergence condition.
+Such uniform stabilization assumptions are standard in classical large-sample theory [andersen1982cox].
+The requirement on the convergence of $s_{k}^{*}/n$ is analogous to the balanced community assumption in lei2015consistency.
+Furthermore, Assumption 2 guarantees that lines characterized in Proposition 3 converge to distinct lines as $n$ diverges, see Lemma S5 in supplementary material for details.
+Assumption 3.
+For $\{\eta_{\Delta,i}^{*}\}$, $i=1,\dots,n$,
+  • (i)
+    There exists $v:(0,1]\rightarrow\mathbb{R}^{+}$ such that for any $\epsilon\in(0,1]$,
+    $v(\epsilon)=\min_{k\in[K]}\liminf_{C\subset\{i,\sigma^{*}(i)=k\},|C|\geq\epsilon n}\sqrt{\frac{1}{|C|}\sum_{i\in C}(\eta_{\Delta,i}^{*}-\frac{1}{|C|}\sum_{i\in C}\eta_{\Delta,i}^{*})^{2}}>0;$
+  • (ii)
+    If $K=2$, there exists $M>0$, such that $\|\bm{\eta}_{\Delta}^{*}\|_{\infty}<M$.
+Assumption 3 (i) is a uniform non-degeneracy condition, which rules out the pathological case where $\eta_{\Delta,i}^{*}$ is nearly constant on any subset of a community, which aligns with the requirement of the identifiability condition in Theorem 1.
+Assumption 3 (ii) requires a uniform upper bound for $|\eta_{\Delta,i}^{*}|$ when there are only two communities.
+Both assumptions can be satisfied if $\eta_{\Delta,i}^{*}$’s within the community come independently from some continuous distribution.
+For example, if $\eta_{\Delta,i}^{*}$ follows a uniform distribution on $[0,1]$, then $v(\epsilon)\simeq\epsilon$.
+Theorem 3.
+Let $\gamma_{n}$ be the smallest nonzero eigenvalue of $\bm{J}\bm{\Theta}^{*}_{\Delta}\bm{J}$.
+Suppose all assumptions in Theorem 2 and Assumptions 2-3 hold, then it holds true with probability $1-n^{-c}$ for positive constant $c$ that
+  • (i)
+    When $K\geq 3$, then there exists a positive constant $C$ that depends on $\{\bar{s}_{k}\}_{k\in[K]}$, $\{\bar{t}_{k}\}_{k\in[K]}$, $\{\bar{u}_{k}\}_{k\in[K]}$ and $K$ such that for sufficiently large $n$,
+    $d(\widehat{\sigma},\sigma^{*})\cdot v(d(\widehat{\sigma},\sigma^{*}))\leq C\sqrt{\frac{\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}_{\Delta}^{*}\|_{F}^{2}}{\gamma^{2}_{n}}+\epsilon_{n}};$ (7)
+  • (ii)
+    When $K=2$, then there exists a positive constant $C$ that depends only on $\{\bar{s}_{k}\}_{k\in[K]}$, $\{\bar{t}_{k}\}_{k\in[K]}$, $\{\bar{u}_{k}\}_{k\in[K]}$, $K$ and $M$ such that for sufficiently large $n$,
+    $\frac{d^{3}(\widehat{\sigma},\sigma^{*})\cdot v^{2}(d(\widehat{\sigma},\sigma^{*}))}{d(\widehat{\sigma},\sigma^{*})+v^{2}(d(\widehat{\sigma},\sigma^{*}))}\leq C\left(\frac{\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}_{\Delta}^{*}\|_{F}^{2}}{\gamma^{2}_{n}}+\epsilon_{n}\right).$ (8)
+Theorem 3 shows that the community detection error depends on the estimation error of $\widehat{\bm{\Theta}}_{\Delta}$, the approximation error $\epsilon_{n}$ and the variability of $\bm{\eta}_{\Delta}^{*}$.
+By Lemma S5, we have $\gamma_{n}\simeq n$, thus the right-hand sides of (7) and (8) converge to zero, and it immediately follows that $d(\widehat{\sigma},\sigma^{*})=o_{p}(1)$ by noting that $v(\cdot)$ is non-decreasing.
 5 Numerical experiments
 In this section, we examine the numerical performance of the proposed SBBM model, and compare it against three popular competitors in the literature, including the signed Laplacian-based method [kunegis2010spectral], the SPONGE algorithm based on signed SBM [cucuringu2021regularized], and the joint estimation method based on the joint inner product model [tang2024population], denoted as SLP, SPONGE and JIM, respectively.
-All methods are evaluated in terms of community detection accuracy, while JIM and SBBM are further compared in terms of probability estimation.
-Note that the community detection error in (4) requires enumeration of all possible permutations, which is computationally infeasible.
-We thus employ the clustering error in wang2010consistent to evaluate the community detection accuracy.
-Particularly, let $\bm{\psi}^{*}=(\psi_{1}^{*},\dots,\psi_{n}^{*})$ and $\widehat{\bm{\psi}}=(\widehat{\psi}_{1},\dots,\widehat{\psi}_{n})$ denote the true and estimated community assignments, then the community detection error of $\widehat{\bm{\psi}}$ can be evaluated as
-$\text{Err}(\widehat{\bm{\psi}},\bm{\psi}^{*})=\frac{2}{n(n-1)}\sum_{i<j}\mathbb{I}\big(\mathbb{I}(\psi_{i}^{*}=\psi_{j}^{*})+\mathbb{I}(\widehat{\psi}_{i}=\widehat{\psi}_{j})=1\big).$
-Further, we define the estimation error of the probability matrix $\bm{P}^{*+}=(p^{*+}_{ij})_{i,j=1}^{n}$ and $\bm{P}^{*-}=(p^{*-}_{ij})_{i,j=1}^{n}$ as
-[…]
+All methods are evaluated in terms of the community detection error defined in (6).
 5.1 Simulated examples
@@
 Edge probabilities are then specified by $\Pr(|A_{ij}|=1)=\sigma(\mu^{*}+\alpha_{i}^{*}+\alpha_{j}^{*}-2\bar{\alpha}^{*}+\bm{x}_{i}^{*\top}\bm{x}_{j}^{*})$ and $\Pr(A_{ij}=1\mid|A_{ij}|=1)=\sigma(v_{i}^{*}v_{j}^{*})$, where $\bar{\alpha}^{*}=n^{-1}\sum_{i=1}^{n}\alpha_{i}^{*}$ and $\sigma(x)=(1+e^{-x})^{-1}$.
-Varying $\mu^{*}\in\{-3.5,-3.0,-2.5\}$ produces strong-balanced signed networks with different sparsity.
+Varying $\mu^{*}\in\{-4.0,-3.5,-3.0\}$ produces strong-balanced signed networks with different sparsity.
 Example 2.
 In this example, for each $i\in[n]$, we first generate $\beta_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, $\gamma_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$, where $\delta_{i}\sim U(0,2)$.
-Then we vary $K\in\{4,6,8\}$ and randomly assign each node to one of the $K$ communities uniformly.
+Then we vary $K\in\{3,4,5\}$ and randomly assign each node to one of the $K$ communities uniformly.
 The matrices $\bm{\Theta}^{*}$ and $\bm{P}^{*}$ are constructed as in (1) and the signed networks are generated accordingly, resulting in weak-balanced signed networks with different numbers of communities.
@@
 In this example, we set $K=4$.
-For each $i\in[n]$, we generate $\beta_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$, $\gamma_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$ with $\sigma^{2}\in\{1,2,3\}$.
+For each $i\in[n]$, we generate $\beta_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$, $\gamma_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$ with $\sigma^{2}\in\{2,3,4\}$.
 We then generate $\delta_{i}\sim U(0,2)$ and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$.
 The signed networks are subsequently generated as in Example 2, resulting in a weak-balanced network with various node heterogeneity.
-We consider $n\in\{500,1000\}$ in all examples.
-The averaged probability estimation errors and community detection errors, along with their standard errors, over 50 independent replications are reported in Tables 1-6.
+We consider $n\in\{500,1000\}$ in all examples. The averaged community detection errors, along with their standard errors, over 50 independent replications are reported in Tables 1-3.
 Table 1: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 1. The best performer in each scenario is boldfaced.
 $n$ $\mu$ SBBM SLP SPONGE JIM
-500 -3.5 0.0910(0.0122) 0.4172(0.0069) 0.3950(0.0106) 0.2073(0.0242)
--3.0 0.0271(0.0014) 0.2822(0.0059) 0.2164(0.0049) 0.1060(0.0189)
--2.5 0.0078(0.0007) 0.1715(0.0050) 0.1169(0.0034) 0.0361(0.0106)
-1000 -3.5 0.0133(0.0007) 0.2016(0.0029) 0.1515(0.0029) 0.0723(0.0162)
--3.0 0.0025(0.0003) 0.1070(0.0023) 0.0711(0.0018) 0.0230(0.0058)
--2.5 0.0004(0.0001) 0.0440(0.0012) 0.0257(0.0009) 0.0161(0.0100)
-Table 2: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 1.
-The better performer in each scenario is boldfaced.
-$n$ $\mu$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
-500 -3.5 0.6085(0.1521) 0.8727(0.2557) 0.5546(0.0126) 0.4829(0.0106)
--3.0 0.3136(0.0014) 0.3964(0.0017) 0.4163(0.0091) 0.3848(0.0094)
--2.5 0.2582(0.0012) 0.3199(0.0011) 0.3187(0.0067) 0.3033(0.0079)
-1000 -3.5 0.2789(0.0008) 0.3527(0.0010) 0.3726(0.0063) 0.3528(0.0069)
--3.0 0.2277(0.0007) 0.2850(0.0007) 0.2819(0.0032) 0.2805(0.0067)
--2.5 0.1888(0.0005) 0.2350(0.0005) 0.2170(0.0048) 0.2140(0.0065)
+500 -4.0 0.1392(0.0076) 0.4481(0.0066) 0.4370(0.0080) 0.2618(0.0204)
+-3.5 0.0909(0.0025) 0.3067(0.0095) 0.2992(0.0160) 0.1536(0.0233)
+-3.0 0.0632(0.0022) 0.1712(0.0047) 0.1239(0.0033) 0.0692(0.0153)
+1000 -4.0 0.0860(0.0024) 0.2357(0.0051) 0.2187(0.0137) 0.1332(0.0201)
+-3.5 0.0578(0.0019) 0.1138(0.0019) 0.0826(0.0017) 0.0481(0.0140)
+-3.0 0.0414(0.0011) 0.0567(0.0013) 0.0369(0.0010) 0.0121(0.0032)
 Table 2: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 2. The best performer in each scenario is boldfaced.
 $n$ $K$ SBBM SLP SPONGE JIM
-500 4 0.0310(0.0041) 0.3201(0.0063) 0.2276(0.0073) -
-6 0.0425(0.0036) 0.3621(0.0049) 0.2692(0.0050) -
-8 0.0508(0.0039) 0.3800(0.0056) 0.2886(0.0051) -
-1000 4 0.0091(0.0006) 0.3128(0.0050) 0.2141(0.0060) -
-6 0.0123(0.0011) 0.3497(0.0037) 0.2600(0.0038) -
-8 0.0152(0.0012) 0.3632(0.0031) 0.2736(0.0032) -
-Table 4: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 2.
-The better performer in each scenario is boldfaced.
-$n$ $K$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
-500 4 0.2265(0.0083) 0.1945(0.0049) 1.0216(0.0093) 0.8595(0.0035)
-6 0.2841(0.0119) 0.1826(0.0038) 1.1016(0.0123) 0.8305(0.0033)
-8 0.3409(0.0128) 0.1819(0.0044) 1.1649(0.0149) 0.8164(0.0033)
-1000 4 0.1546(0.0035) 0.1344(0.0024) 1.0343(0.0096) 0.8587(0.0026)
-6 0.1788(0.0049) 0.1225(0.0020) 1.1188(0.0130) 0.8290(0.0025)
-8 0.1993(0.0053) 0.1168(0.0016) 1.1869(0.0159) 0.8151(0.0024)
+500 3 0.0379(0.0024) 0.0857(0.0109) 0.1296(0.0063) -
+4 0.0687(0.0046) 0.3170(0.0063) 0.2302(0.0072) -
+5 0.1215(0.0098) 0.3799(0.0057) 0.2938(0.0077) -
+1000 3 0.0174(0.0009) 0.0363(0.0080) 0.0955(0.0062) -
+4 0.0312(0.0014) 0.3114(0.0051) 0.2194(0.0080) -
+5 0.0591(0.0049) 0.3750(0.0038) 0.2910(0.0056) -
 Table 3: Averaged community detection errors and their standard errors of all methods over 50 independent replications in Example 3. The best performer in each scenario is boldfaced.
 $n$ $\sigma^{2}$ SBBM SLP SPONGE JIM
-500 1 0.1858(0.0063) 0.2355(0.0071) 0.2603(0.0229) -
-2 0.1418(0.0082) 0.2802(0.0073) 0.1675(0.0089) -
-3 0.0784(0.0073) 0.2951(0.0051) 0.1662(0.0080) -
-1000 1 0.0845(0.0032) 0.1855(0.0074) 0.0970(0.0077) -
-2 0.0465(0.0015) 0.2640(0.0044) 0.0882(0.0049) -
-3 0.0226(0.0010) 0.2749(0.0044) 0.1208(0.0080) -
-Table 6: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 3.
-The better performer in each scenario is boldfaced.
-$n$ $\sigma^{2}$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
-500 1 0.4052(0.0063) 0.8484(0.0184) 1.6358(0.0486) 3.8576(0.0938)
-2 0.3283(0.0089) 0.4778(0.0203) 1.2363(0.0253) 1.5960(0.0327)
-3 0.2717(0.0092) 0.3165(0.0103) 1.1129(0.0168) 1.1200(0.0138)
-1000 1 0.2725(0.0026) 0.5233(0.0055) 1.6849(0.0516) 3.7927(0.0585)
-2 0.2084(0.0021) 0.2893(0.0035) 1.2639(0.0263) 1.5614(0.0211)
-3 0.1751(0.0021) 0.2085(0.0024) 1.1328(0.0175) 1.1076(0.0092)
-It is evident that SBBM delivers the best performance in terms of both probability estimation and community detection in almost all examples, except that its probability estimation accuracy is comparable to JIM in Example 1 with strong-balanced signed networks.
+500 2 0.1396(0.0038) 0.2735(0.0058) 0.1781(0.0095) -
+3 0.1036(0.0038) 0.2850(0.0057) 0.1796(0.0114) -
+4 0.0776(0.0031) 0.2851(0.0057) 0.1689(0.0072) -
+1000 2 0.0850(0.0053) 0.2616(0.0035) 0.1187(0.0142) -
+3 0.0382(0.0039) 0.2780(0.0046) 0.1140(0.0088) -
+4 0.0275(0.0012) 0.2754(0.0047) 0.1283(0.0079) -
+It is evident that SBBM delivers the best performance in terms of community detection in almost all examples, except that it is comparable to JIM in Example 1 with strong-balanced signed networks.
 Note that JIM is not applicable to detect communities in weak-balanced signed networks with $K\geq 2$, and thus the blanks in Tables 3 and 5.
-Its probability estimation accuracy also becomes substantially less competitive than SBBM in Examples 2 and 3 with weak-balanced signed networks.
-Furthermore, across all examples, the community detection errors and probability estimation errors of SBBM decrease significantly with $n$, confirming its asymptotic consistencies in Theorems 2 and 3.
+Furthermore, across all examples, the community detection errors of SBBM decrease significantly with $n$, confirming its asymptotic consistencies in Theorems 3.
 In contrast, the performance of SLP and SPONGE improves slowly with $n$ in Examples 2 and 3 with node heterogeneity, which is possibly due to the fact that node heterogeneity is largely ignored by SLP and SPONGE.
@@
 5.2 International relation network
-We now turn to apply SBBM to analyze an international relation network constructed from two sources, the BACI trade database (CEPII, version 2025-01) [CEPII:2010-23], which harmonizes annual bilateral trade flows, and the Global Sanctions Data Base R4 [felbermayr2020global; yalcin2025global] that catalogs economic sanctions between economies.
-For each pair of economies, if economic sanctions have ever existed between them during 2015-2023, a negative edge is assigned; otherwise, a positive edge is assigned if their bilateral trade volume ranks in the top $1/30$.
-This pre-processing leads to a signed network with 164 economies, 580 positive edges, and 589 negative edges.
-It contains 2,756 strong-balanced triads, 3,559 weak-balanced triads, and 847 unbalanced triads.
-It is evident that weak-balanced triads occur more than four times as frequently as unbalanced ones, in sharp contrast to the expected ratio 5:3 under random assignment.
+We now apply SBBM to analyze an international relation network constructed from two sources, the BACI trade database (CEPII, version 2025-01) [CEPII:2010-23], which harmonizes annual bilateral trade flows, and the Global Sanctions Data Base R4 [felbermayr2020global; yalcin2025global] that catalogs economic sanctions between economies.
+For each pair of economies, if economic sanctions have ever existed between them during 2022–2024, a negative edge is assigned; otherwise, a positive edge is assigned if their bilateral trade volume ranks in the top 5 for either economy.
+This pre-processing leads to a signed network with 230 economies, 947 positive edges, and 1,476 negative edges.
+It contains 6,451 strong-balanced triads, 10209 weak-balanced triads, and 605 unbalanced triads.
+It is evident that weak-balanced triads occur more than sixteen times as frequently as unbalanced ones, in sharp contrast to the expected ratio 5:3 under random assignment.
 This substantial deviation from randomness indicates a pronounced weak balance structure in the international relation network.
-The community structure in the international relation network is then estimated via SBBM, SLP, SPONGE and JIM, respectively.
-Also, following the suggestion in zhang2022signed, the number of communities $K=4$ is determined by optimizing the Bayesian information criterion [saldana2017many].
-For comparison, we compute the signed modularity gomez2009analysis of estimated communities from different methods.
+The community structure in the international relation network is then estimated via SBBM, SLP, and SPONGE with $K=3$, which is determined by the Gap statistic [tibshirani2001estimating; zhao2011community], and via JIM with $K=2$.
+For comparison, we compute the signed modularity [gomez2009analysis] of estimated communities from different methods.
 Particularly, for each $i\in[n]$, let $k^{+}_{i}=\sum_{j\in[n]}a^{+}_{ij}$, $k^{-}_{i}=\sum_{j\in[n]}a^{-}_{ij}$, $m^{+}=\frac{1}{2}\sum_{i,j\in[n]}a^{+}_{ij}$, and $m^{-}=\frac{1}{2}\sum_{i,j\in[n]}a^{-}_{ij}$.
-The signed modularity of a given community assignment $\{\psi_{i}\}_{i\in[n]}$ is defined as
-$Q_{\text{signed}}=\frac{1}{2m^{+}}\sum_{i,j\in[n]}\left(a^{+}_{ij}-\frac{k_{i}^{+}k_{j}^{+}}{2m^{+}}\right)\delta(\psi_{i},\psi_{j})-\frac{1}{2m^{-}}\sum_{i,j\in[n]}\left(a^{-}_{ij}-\frac{k_{i}^{-}k_{j}^{-}}{2m^{-}}\right)\delta(\psi_{i},\psi_{j}),$
-where $\delta(\psi_{i},\psi_{j})=1$ if $\psi_{i}=\psi_{j}$ and 0 otherwise.
+The signed modularity of a given community assignment $\{\sigma(i)\}_{i\in[n]}$ is defined as
+$Q_{\text{signed}}=\frac{1}{2m^{+}}\sum_{i,j\in[n]}\left(a^{+}_{ij}-\frac{k_{i}^{+}k_{j}^{+}}{2m^{+}}\right)\delta(\sigma(i),\sigma(j))-\frac{1}{2m^{-}}\sum_{i,j\in[n]}\left(a^{-}_{ij}-\frac{k_{i}^{-}k_{j}^{-}}{2m^{-}}\right)\delta(\sigma(i),\sigma(j)),$
+where $\delta(\sigma(i),\sigma(j))=1$ if $\sigma(i)=\sigma(j)$ and 0 otherwise.
 Note that $Q_{\text{signed}}$ takes values in [2,2][-2,2], and a larger value indicates more positive edges within communities and more negative edges between communities.
@@
 SBBM SLP SPONGE JIM
-$Q_{\text{signed}}$ 0.4266 0.1943 0.1296 -0.0032
-Figure 2 further displays the heatmap of the international relation network, where nodes are reordered according to $\widehat{\bm{Z}}$ from SBBM, and positive and negative edges are dyed into red or blue.
-It is clear that most positive edges occur within the same community, while negative edges occur across different communities, and a particularly high concentration of positive edges can be observed in the largest community.
+$Q_{\text{signed}}$ 0.7012 0.5099 0.4163 3.1717e-5
+Figure 2 further displays the heatmap of the international relation network, where nodes are reordered according to $\widehat{\sigma}$ from SBBM, and positive and negative edges are dyed into red or blue.
+It is clear that most positive edges occur within the same community, while negative edges occur across different communities, and a particularly high concentration of negative edges can be observed between the first and second communities.
 Figure 2: Heatmap of the international relation network.
 We also plot the estimated communities of SBBM on a world map in Figure 3.
-The sizes of the four estimated communities are 106, 29, 24, and 5, respectively.
-The largest one is centered on China and mainly consists of developing economies in Asia, Africa and Latin America.
-The second community is centered on the United States and mainly consists of developed economies in Europe and North America.
-The third community mainly consists of Afghanistan, Belarus, Iran, Syria, North Korea, Russia, Sudan and Lebanon, and the fourth community consists of Bosnia and Herzegovina, Cyprus, Eritrea, Moldova, and Ukraine.
-The economies in these two communities are conflict-affected or frequently subject to sanctions from other economies, and more interestingly, there are even stronger sanction ties between them.
-We further scrutinize the patterns of positive and negative edges via the estimated parameters.
-It is noteworthy that both China and the United States have high $\widehat{\beta}^{+}$, indicating their leading roles in forming strong within-community trading relations.
-On the other hand, both the United States and the United Kingdom have high $\widehat{\gamma}^{-}$, indicating their leading roles in imposing sanctions.
-Figure 3: The four estimated communities by SBBM on the world map, where economics in different communities are dyed with different colors.
+The sizes of the three estimated communities are 90, 87, 53, respectively.
+The first community is centered on the United States and the European Union, encompassing most developed economies and their regional partners in the Americas.
+The second community is anchored by China and Russia, aggregating major emerging economies such as India, Saudi Arabia that typically do not align with Western-led sanctions.
+The third community forms a distinct cluster in the Asia-Pacific region, comprising Japan, South Korea, Australia, and members of the Association of Southeast Asian Nations.
+This tripartite structure aligns with recent findings on geoeconomic fragmentation [gopinath2025changing], confirming the divergence into U.S.- and China-centric spheres.
+Crucially, the result identifies an intermediate Pacific economic community that maintains regional integration despite divergent security alliances.
+Figure 3: The three estimated communities by SBBM on the world map, where economics in different communities are dyed with different colors.
 6 Summary
@@
 It introduces four sets of parameters that distinguish each node’s connection tendencies within its community and across different communities, which greatly facilitates community detection in signed networks with balance structures and node heterogeneity.
-We establish the identifiability conditions for the signed block $\beta$-model, and develop an efficient alternative updating algorithm for parameter estimation and community detection.
-Its asymptotic properties are also established in terms of both probability estimation and community detection, which are further supported by numerical experiments on both synthetic and real-world experiments.
+We establish the identifiability conditions for the signed block $\beta$-model, and develop a two-step algorithm for parameter estimation and community detection.
+Its asymptotic properties are also established in terms of both parameter estimation and community detection, which are further supported by numerical experiments on both synthetic and real-world experiments.
 Note that statistical inference about the signed block $\beta$-model is also of interest but possibly hindered by the discrete feature of the community structure, and we leave it to future investigation.